### v2

# From graph to Riesz continuity

Abstract
We show that every graph continuous family of unbounded operators in a Hilbert space becomes Riesz continuous after multiplication by an appropriate family of unitary operators. This result provides a simple definition of the index for graph continuous families of Fredholm operators and leads to two corollaries for operators with compact resolvents: (1) the identity map between the spaces of such operators equipped with the Riesz and the graph topology is a homotopy equivalence; (2) every graph continuous family of such operators acting between fibers of Hilbert bundles becomes Riesz continuous in appropriate trivializations of the bundles.
For self-adjoint operators, multiplication by unitaries should be replaced by conjugation. In general, a graph continuous family of self-adjoint operators with compact resolvents cannot be made Riesz continuous by an appropriate conjugation. We obtain a partial analogue of the “trivialization” result above for self-adjoint operators and describe obstructions to existence of such a trivialization arising in the general case. This part of results is closely related to the recent work [$I_{2}$] of N. Ivanov.
footnotetext: Department of Mathematics, Technion – Israel Institute of Technology.footnotetext: This work was partially supported by ISF grants no. 431/20 and 844/19.
1 Introduction
Let $H$ and $H^{\prime}$ be separable complex Hilbert spaces of infinite dimension. We denote by $\mathcal{B}(H,H^{\prime})$ the space of bounded linear operators $H\to H^{\prime}$ with the norm topology; by $\mathcal{K}(H,H^{\prime})$ the subspace of $\mathcal{B}(H,H^{\prime})$ consisting of compact operators, by $\mathcal{U}(H)$ and $\mathcal{P}(H)$ the subspaces of $\mathcal{B}(H)=\mathcal{B}(H,H)$ consisting of unitary operators and projections respectively (by projections we always mean orthogonal projections).
Regular operators.   An unbounded operator $A$ from $H$ to $H^{\prime}$ is a linear operator defined on a subspace $\operatorname{dom}(A)$ of $H$ and taking values in $H^{\prime}$. Such an operator $A$ is called closed if its graph is closed in $H\oplus H^{\prime}$, and densely defined if its domain $\operatorname{dom}(A)$ is dense in $H$. It is called regular if it is closed and densely defined. Let $\mathcal{R}(H,H^{\prime})$ denote the set of all regular operators $H\to H^{\prime}$ and $\mathcal{R}^{\mathrm{sa}}(H)\subset\mathcal{R}(H)=\mathcal{R}(H,H)$ denote the subset of self-adjoint operators.
Two topologies on regular operators.   The set $\mathcal{R}(H,H^{\prime})$ of regular operators has two important natural topologies. The Riesz topology on $\mathcal{R}(H,H^{\prime})$ is induced by the inclusion
$\bm{\chi}\colon\mathcal{R}(H,H^{\prime})\hookrightarrow\mathcal{B}(H,H^{\prime}),\quad\bm{\chi}(A)=A(1+A^{*}A)^{-1/2}$
from the norm topology on the space of bounded operators. Here $\bm{\chi}$ is the so called bounded transform map. The graph topology on $\mathcal{R}(H,H^{\prime})$ is induced by the inclusion
$\mathbf{p}\colon\mathcal{R}(H,H^{\prime})\hookrightarrow\mathcal{P}(H\oplus H^{\prime})$
from the norm topology on the space of projections, where $\mathbf{p}$ is the map taking a regular operator to the orthogonal projection onto its graph. Let ${{}^{r}\mathcal{R}}(H,H^{\prime})$, resp. ${{}^{g}\mathcal{R}}(H,H^{\prime})$ denote the space of regular operators $H\to H^{\prime}$ equipped with the Riesz, resp. graph topology.
The Riesz topology is strictly stronger than the graph topology. In other words, the identity map ${{}^{r}\mathcal{R}}\to{{}^{g}\mathcal{R}}$ is continuous, while the identity map ${{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ is not. On the subset $\mathcal{B}(H,H^{\prime})$ of bounded operators these two topologies coincide.
From graph to Riesz continuity.   Riesz continuous families of operators are much easier to deal with than mere graph continuous families. This motivates the approach that we are developing in this paper.
Suppose that one is given a graph continuous family of regular operators. One can try to modify it, in an appropriate way, so that the resulting family becomes Riesz continuous, then investigate the modified family and transfer results of this investigation to the original family. By an “appropriate way” we mean that such a transfer of results is possible, that is, some important properties of the original family may be recovered from the corresponding properties of the modified family. Of course, this description is quite vague; we will make it more precise below.
Let us illustrate our approach by an example of a family consisting of invertible operators. Recall that an operator $A\in\mathcal{R}(H)$ is called invertible if it is bijective as a map $\operatorname{dom}(A)\to H$. Such operator has a unique polar decomposition $A=u\cdot|A|$, where $u\in\mathcal{U}(H)$ is a unitary operator and $|A|=\sqrt{A^{*}A}$ is a positive operator. If $\mathcal{A}\colon X\to\mathcal{R}(H)$ is graph continuous, then the map $x\mapsto|\mathcal{A}_{x}|$ is Riesz continuous. If additionally all operators $\mathcal{A}_{x}$ are invertible, then the map $x\mapsto u_{x}=\mathcal{A}_{x}|\mathcal{A}_{x}|^{-1}$ from $X$ to ${{}^{s}{\mathcal{U}}}(H)$, the unitary group equipped with the strong operator topology, is continuous. Taking $v_{x}=u_{x}^{*}$, we obtain a continuous map $v\colon X\to{{}^{s}{\mathcal{U}}}(H)$ such that the map $x\mapsto v_{x}\mathcal{A}_{x}=|\mathcal{A}_{x}|$ is Riesz continuous. The original family $\mathcal{A}$ can be recovered from the modified family $|\mathcal{A}|$ by the left multiplication by $u_{x}$.
This construction cannot be extended continuously to general non-invertible families because the polar decomposition of a general non-invertible operator involves a partial isometry which is not unitary. Moreover, this partial isometry does not depend continuously on the operator. However, one can still modify a family to a Riesz continuous one using the multiplication by an appropriate family of unitary operators, as we show in Theorem A below (we will multiply on the right, but it really does not matter: one can always pass to adjoint operators, replacing the left multiplication with the right one). In contrast with the construction above, the modified family no longer consists of positive operators, nor is it determined by the original family in any canonical way. But the modified family still carries an important information about the original one.
Theorem A.
There exists a map $v\colon\mathcal{R}(H,H^{\prime})\to\mathcal{U}(H)$, which is both Riesz-to-norm and graph-to-strong continuous, such that the induced map
$\Phi\colon\mathcal{R}(H,H^{\prime})\to\mathcal{R}(H,H^{\prime}),\quad A\mapsto A\cdot v(A)$
is graph-to-Riesz continuous.
[…]
$K^{0}$ index for graph continuous families.   Let $\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}$ be a graph continuous family and $\Phi\colon{{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ be a map provided by Theorem A. Then $\bm{\chi}\circ\Phi\circ\mathcal{A}=\varphi\circ\mathcal{A}$ is a norm continuous family of bounded operators. Suppose that all $\mathcal{A}_{x}$ are Fredholm. Then the operators $\varphi(\mathcal{A}_{x})$ are also Fredholm, and one can define the index of $\mathcal{A}$ as the index of this norm continuous family of bounded Fredholm operators:
$\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\varphi\circ\mathcal{A})=\operatorname{ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})\in[X,\mathcal{B}_{F}]=K^{0}(X).$
This definition of the index is discussed in the last part of Section 3; it is justified by Theorem B below.
Equivalently, the index of $\mathcal{A}$ may be defined as the index of the family of projections $\Gamma_{x}\to H^{\prime}$ after a trivialization of the Hilbert bundle $\Gamma$ formed by the graphs $\Gamma_{x}$ of operators $\mathcal{A}_{x}$. (This trivialization should be chosen in the category of bundles with the structure group $\mathcal{U}$, not ${{}^{s}{\mathcal{U}}}$.) Note that such a trivialization is not canonical; it is defined only up to a continuous map $X\to\mathcal{U}(H)$. However, change of a trivialization does not affect the homotopy type of the corresponding family of bounded Fredholm operators, since the unitary group $\mathcal{U}(H)$ is contractible.
Searching for such a definition of the index was part of the motivation behind Theorem A. An alternative approach to the index of graph continuous families of Fredholm operators is provided by the previous paper [$P_{3}$] of the author. Another approach, for compact base spaces and in a more general framework of Hilbert modules over C*-algebras, is provided by results of M. Joachim [J]. Recently N. Ivanov [$I_{1}$, $I_{2}$] developed a new approach to the index of families, which works under much weaker continuity assumptions than graph continuity. We show in the end of Section 3 that for graph continuous families of Fredholm operators our definition of the index given above coincides with Ivanov’s index.
Homotopy equivalences.   A regular operator $A$ is said to have compact resolvents***If the resolvent set $\operatorname{Res}(A)$ of $A$ is non-empty (for example, if $A$ is self-adjoint), then this definition agrees with the usual sense of the words “compact resolvent”: $(A-\lambda)^{-1}$ is compact for some (and then every) $\lambda\in\operatorname{Res}(A)$. However, a regular operator may have an empty resolvent set; the definition above covers such operators as well. if both $(1+A^{*}A)^{-1}$ and $(1+AA^{*})^{-1}$ are compact operators. Let $\mathcal{R}_{F}(H)$ and $\mathcal{R}_{K}(H)$ be the subsets of $\mathcal{R}(H)$ consisting of Fredholm operators and operators with compact resolvents, respectively. These subsets are invariant under the multiplication by unitary operators, so $\Phi(\mathcal{R}_{K})\subset\mathcal{R}_{K}$ and $\Phi(\mathcal{R}_{F})\subset\mathcal{R}_{F}$.
It was shown in the previous paper of the author that the identity maps
${{}^{r}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}\;\;\text{and}\;\;{{}^{r}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$
are homotopy equivalences. See [$P_{3}$, Theorem A]. However, the proof in [$P_{3}$] provides no intuition on how a homotopy inverse map might look like. The following theorem reveals a possible form of homotopy inverse maps and provides an alternative proof of homotopy equivalence of the identity map ${{}^{r}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$.
Theorem B.
Let $\Phi\colon{{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ be a map provided by Theorem A. Then the restriction of $\Phi$ to the subspace of operators with compact resolvents is homotopy inverse to the identity map ${{}^{r}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$. Similarly, the restriction of $\Phi$ to the subspace of Fredholm operators is homotopy inverse to the identity map ${{}^{r}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$.
Let $J$ be a subset of $\mathcal{K}(H)$ invariant under the conjugation by unitary operators. Then one can replace $\mathcal{R}_{K}$ in the statement of Theorem B with the set
$\mathcal{R}_{J}(H)=\left\{A\in\mathcal{R}(H)\,|\,(1+A^{*}A)^{-1},(1+AA^{*})^{-1}\in J\right\}.$
(In particular, $J$ can be the subset of trace class operators, or the subset of Hilbert–Schmidt operators.) Such a set $\mathcal{R}_{J}$ is invariant under the multiplication by unitary operators, so both $\Phi$ and homotopies used in the proof of Theorem B preserve $\mathcal{R}_{J}$. This provides the following generalization of the first part of Theorem B.
Theorem B’. Let $J$ be a subset of $\mathcal{K}(H)$ invariant under the conjugation by unitary operators. Then the identity map ${{}^{r}\mathcal{R}}_{J}(H)\to{{}^{g}\mathcal{R}}_{J}(H)$ is a homotopy equivalence, with a homotopy inverse map given by $\Phi\colon{{}^{g}\mathcal{R}}_{J}\to{{}^{r}\mathcal{R}}_{J}$.
Hilbert bundles.   Let $\mathcal{H}$ and $\mathcal{H}^{\prime}$ be locally trivial Hilbert bundles over a topological space $X$ with a fiber $H$. The structure group of a general Hilbert bundle is the unitary group ${{}^{s}{\mathcal{U}}}(H)$ equipped with the strong operator topology. As was shown by Dixmier and Douady, every such bundle over a paracompact base space $X$ is trivial, that is, isomorphic to the trivial Hilbert bundle $X\times H\to X$. See [DD], Theorem 1. More generally, every numerable Hilbert bundle is trivial. However, automorphisms of the trivial Hilbert bundle $X\times H\to X$ are given by continuous maps $X\to{{}^{s}{\mathcal{U}}}(H)$ and preserve neither norm nor Riesz continuity of operator families, since the action of ${{}^{s}{\mathcal{U}}}(H)$ on $\mathcal{B}(H)$ by conjugation, $(u,a)\mapsto uau^{-1}$, is not continuous. Therefore, the notions of a norm continuous family of bounded operators or of a Riesz or graph continuous family of regular operators are not defined for a general Hilbert bundle.
Nevertheless, the actions of ${{}^{s}{\mathcal{U}}}(H)$ and ${{}^{s}{\mathcal{U}}}(H^{\prime})$ on the space $\mathcal{K}(H,H^{\prime})$ of compact operators are continuous, and thus the notion of a norm continuous family of compact operators $k_{x}\colon\mathcal{H}_{x}\to\mathcal{H}^{\prime}_{x}$ is well defined. By the same reason, the notion of a graph continuous family of regular operators $\mathcal{A}_{x}\colon\mathcal{H}_{x}\to\mathcal{H}^{\prime}_{x}$ with compact resolvents is well defined. Applying Theorem A to such a family, we obtain the following result.
Theorem C.
Let $\mathcal{H}$ and $\mathcal{H}^{\prime}$ be numerable Hilbert bundles over a topological space $X$. Let $\mathcal{A}$ be a graph continuous family of regular operators $\mathcal{A}_{x}\in\mathcal{R}_{K}(\mathcal{H}_{x},\mathcal{H}^{\prime}_{x})$ with compact resolvents. Suppose that a global trivialization of one of these Hilbert bundles, $\mathcal{H}\cong X\times H$ or $\mathcal{H}^{\prime}\cong X\times H^{\prime}$, is fixed. Then there is a global trivialization of the second bundle taking $\mathcal{A}$ to a Riesz continuous map $X\to\mathcal{R}_{K}(H,H^{\prime})$.
We prove Theorems B and C in Section 3.
Self-adjoint operators.   For self-adjoint operators, multiplication $A\mapsto Au$ or $A\mapsto uA$ should be replaced by the conjugation $A\mapsto uAu^{-1}$. Instead of independent trivializations of two Hilbert bundles $\mathcal{H}$ and $\mathcal{H}^{\prime}$, as in Theorem C, one should look for a one trivialization of a given Hilbert bundle $\mathcal{H}$. This narrows down available options drastically. In particular, Theorem A has no direct analogue for self-adjoint operators, even invertible ones. A graph continuous family of self-adjoint operators with compact resolvents may fail to be Riesz continuous in any trivialization of $\mathcal{H}$, as we will explain below. The next theorem shows that the obstruction to the existence of such a trivialization has purely local nature.
From local to global trivializations.   Let $\mathcal{A}$ be a family of regular operators $\mathcal{A}_{x}\in\mathcal{R}(\mathcal{H}_{x})$ acting on fibers of a Hilbert bundle $\mathcal{H}$. Following Ivanov [$I_{2}$], we say that a trivialization $\tau\colon\left.\mathcal{H}\right|_{Y}\to Y\times H$ of $\mathcal{H}$ over a subset $Y\subset X$ is adapted to $\mathcal{A}$ if it takes $\mathcal{A}$ to a Riesz continuous map $\tau_{*}\mathcal{A}\colon Y\to\mathcal{R}(H)$. We say that $\mathcal{A}$ is Riesz quasi-continuous if, for every $x\in X$, it admits an adapted trivialization over a neighborhood of $x$.
Theorem D.
Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a paracompact space $X$, and let $\mathcal{A}$ be a family of self-adjoint operators with compact resolvents acting on fibers of $\mathcal{H}$. Suppose that $\mathcal{A}$ is Riesz quasi-continuous (that is, admits local adapted trivializations). Then there exists a global trivialization $\tau\colon\mathcal{H}\to X\times H$ adapted to $\mathcal{A}$. Moreover, every adapted trivialization of $\mathcal{H}$ over a closed subset $Y$ of $X$ can be extended to a global adapted trivialization.
Applying the last statement of the theorem to the lifting of $\mathcal{A}$ to $X\times[0,1]$, we see that every two adapted trivializations are homotopic and thus the homotopy class of a Riesz continuous map $\tau_{*}\mathcal{A}\colon X\to\mathcal{R}_{K}^{\mathrm{sa}}$ does not depend on the choice of an adapted trivialization $\tau$. For a family $\mathcal{A}$ satisfying assumptions of the theorem, this allows to define the $K^{1}$ index of $\mathcal{A}$ as the index of the norm continuous family $\bm{\chi}(\tau_{*}\mathcal{A})$ of bounded Fredholm self-adjoint operators.
The main case of Theorem D is the case of families consisting of operators which are neither essentially positive nor essentially negative. In this case a similar trivialization theorem was recently proved by different methods and under much weaker continuity assumptions by N. Ivanov. See [$I_{2}$], Theorems 4.6 and 4.7. In combination with Lemma 5.1 from his paper [$I_{3}$], his trivialization theorem implies Theorem D for families consisting of operators which are neither essentially positive nor essentially negative.
Reading Ivanov’s paper [$I_{2}$] led the author to realizing that a part of the original proof of Theorem D can be separated from the rest and has an independent interest. In more details, Ivanov mentioned in the introduction to the first version of [$I_{2}$] that there is a notion of a polarization of a Hilbert bundle standing behind some proofs in [$I_{2}$] (he worked out the details in the current version of [$I_{2}$]). In Section 4 we introduce the notion of a polarization of a Hilbert bundle related to Ivanov’s polarizations in the same way as compact operators are related to finite rank ones, and prove an analogue of his trivialization theorem. See Theorem 4.1.
From graph to Riesz continuity.   Let $\mathcal{A}$ be a graph continuous family of self-adjoint operators with compact resolvents acting on fibers of a Hilbert bundle $\mathcal{H}$, as above. The first obstruction to existence of an adapted trivialization is caused by the fact that $\mathcal{R}^{\mathrm{sa}}_{K}$ is path connected in the graph topology, but has three connected components in the Riesz topology. Two points from different components of ${{}^{r}\mathcal{R}}_{K}^{\mathrm{sa}}$ (for example, an essentially positive and an essentially negative operator) may be connected by a graph continuous path in $\mathcal{R}_{K}^{\mathrm{sa}}(H)$, but every trivialization of the trivial Hilbert bundle $[0,1]\times H$ will leave this path Riesz discontinuous.
When all the operators $\mathcal{A}_{x}$ are contained in a fixed connected component of ${{}^{r}\mathcal{R}}_{K}^{\mathrm{sa}}$, there is another kind of local obstruction which we now describe. Let $\lambda\in\mathbb{R}$ be chosen outside of the spectrum of $\mathcal{A}_{x_{0}}$, $x_{0}\in X$. Then the spectral projections $p_{x}=\operatorname{\mathbb{1}}_{[\lambda,+\infty)}(\mathcal{A}_{x})$ (here $\operatorname{\mathbb{1}}_{S}$ denotes the characteristic function of a subset $S\subset\mathbb{R}$) give rise to the Hilbert bundles $\mathcal{H}^{+}$ and $\mathcal{H}^{-}$, with fibers $\mathcal{H}^{+}_{x}=\operatorname{Ran}(p_{x})$ and $\mathcal{H}^{-}_{x}=\operatorname{Ker}(p_{x})$, over some neighborhood of $x_{0}$. A local adapted trivialization of $\mathcal{H}$ near $x_{0}$ exists if and only if both $\mathcal{H}^{+}$ and $\mathcal{H}^{-}$ are locally trivial near $x_{0}$. See Proposition 5.3.
Suppose first that all the operators $\mathcal{A}_{x}$ are essentially negative. Since the norm and strong topologies coincide on the space of projections of a fixed finite rank, an obstruction to existence of local adapted trivializations near $x_{0}$ arises when the rank of $p_{x}$ is not locally constant near $x_{0}$. In other words, there are eigenvalues of $\mathcal{A}_{x}$ going to +$\infty$+ ∞ when $x\to x_{0}$. If this is not the case and the operators $\mathcal{A}_{x}$ are uniformly bounded from above in a neighborhood of each point $x_{0}\in X$, then $\mathcal{A}$ is Riesz continuous in every trivialization of $\mathcal{H}$. The situation with essentially positive operators is completely the same, but +$\infty$+ ∞ and bounds from above should be replaced by $\infty$- ∞ and bounds from below. See, e.g., Rellich’s example [Ka, Example V-4.14].
Suppose now that all $\mathcal{A}_{x}$ are contained in the third connected component of ${{}^{r}\mathcal{R}}_{K}^{\mathrm{sa}}(H)$, ${{}^{r}\mathcal{R}}_{K}^{\star}(H)$, which consists of operators that are neither essentially positive nor essentially negative. Then both Hilbert bundles $\mathcal{H}^{+}$ and $\mathcal{H}^{-}$ have infinite rank. By a result of Dixmier and Douady, every separable Hilbert bundle of infinite rank over a paracompact base space of finite covering dimension is trivial. Recall that $X$ is said to have finite covering dimension if there is an $n\in\mathbb{N}$ such that every finite open covering of $X$ has a finite open refinement of order $\leqslant n$.
Let us say that a space has locally finite covering dimension if each point has a neighborhood of finite covering dimension. Using the aforementioned result of Dixmier and Douady, we prove a partial analogue of Theorem C for self-adjoint operators.
Theorem E.
Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a metrizable space $X$ of locally finite covering dimension. Let $\mathcal{A}$ be a graph continuous family of self-adjoint operators $\mathcal{A}_{x}\in\mathcal{R}^{\star}_{K}(\mathcal{H}_{x})$, with compact resolvents, that are neither essentially positive nor essentially negative. Then $\mathcal{A}$ becomes Riesz continuous in some global trivialization of $\mathcal{H}$. Moreover, every adapted trivialization of $\mathcal{H}$ over a closed subset $Y$ of $X$ can be extended to a global adapted trivialization.
Since the group ${{}^{s}{\mathcal{U}}}(H)$ is contractible, Theorem E implies that the identity map ${{}^{r}\mathcal{R}}^{\star}_{K}\to{{}^{g}\mathcal{R}}^{\star}_{K}$ is a weak homotopy equivalence. The stronger result, of a homotopy equivalence of this map, is proven in the previous paper of the author by a different method. See [$P_{3}$], Theorems B and C.
Theorem E does not hold without additional assumptions for general metrizable base spaces, even compact, as we demonstrate by Example 5.5.
Elliptic operators on closed manifolds.   Let us describe an important class of situations where Theorems C and D may be useful. Let $\mathcal{M}\to X$ be a smooth fibration whose fibers $\mathcal{M}_{x}$ are smooth closed manifolds, and let $\mathcal{E}$, $\mathcal{E}^{\prime}$ be smooth vector bundles over $\mathcal{M}$. Let $\mathcal{D}_{x}$ be a continuous (in an appropriate sense) family of elliptic pseudodifferential operators on $\mathcal{M}_{x}$, of positive order, acting from fibers of $\mathcal{E}$ to fibers of $\mathcal{E}^{\prime}$. A local trivialization of $\mathcal{M}$ turns $\mathcal{D}$ to a family of elliptic operators acting between fibers of fixed vector bundles $E$ and $E^{\prime}$ over a fixed closed manifold $M$. The bounded transform takes this new family to a continuous family of zeroth order elliptic operators, which is norm continuous as a family of bounded operators $L^{2}(M;E)\to L^{2}(M;E^{\prime})$. Hence the family of unbounded operators $\mathcal{D}_{x}$ is Riesz continuous in this local trivialization. However, the fibration $\mathcal{M}\to X$ is in general non-trivial, and thus no such “geometric” global trivialization exists in the general case. The structure group of the corresponding Hilbert bundles $\mathcal{H}$ and $\mathcal{H}^{\prime}$, where $\mathcal{H}_{x}=L^{2}(\mathcal{M}_{x};\mathcal{E}_{x})$ and $\mathcal{H}^{\prime}_{x}=L^{2}(\mathcal{M}_{x};\mathcal{E}^{\prime}_{x})$, has the topology coarser than the norm topology, so Kuiper’s theorem cannot be used in order to trivialize $\mathcal{H}$ and $\mathcal{H}^{\prime}$. Dixmier-Douady theorem, on the other hand, can be used, but trivializations it provides do not take $\mathcal{D}$ to a Riesz continuous family of operators.
Theorem C closes this gap, showing that such a family $\mathcal{D}$ can be made Riesz continuous by appropriate global trivializations of the Hilbert bundles $\mathcal{H}$ and $\mathcal{H}^{\prime}$, though at the cost of losing the differential and geometric structure. Theorem D solves the same problem for families of self-adjoint elliptic operators acting on fibers of $\mathcal{E}$.
The author arrived at Theorem D in an attempt to understand a step in the proof of a result of Melrose and Piazza. See [$MP_{1}$], Proposition 1. Melrose and Piazza deal with a family of self-adjoint differential operators of first order over a smooth fibration $\mathcal{M}\to X$, as above, and show that if the $K^{1}$-index of the family vanishes, then the family admits a spectral section. In their proof, they claim that the $L^{2}$ spaces of the fibers, $L^{2}(\mathcal{M}_{x};\mathcal{E}_{x})$, can be identified so that all operators act on a fixed Hilbert space $H$. Unfortunately, they did not explain how such an identification can be obtained. The author was not able to find a simple explanation of this step. Our Theorem D, as well as recent results [$I_{2}$, Theorems 4.5 or 4.6] or [$I_{3}$] of N. Ivanov, allows to overcome this difficulty. Similarly, Theorem C allows to pass to a fixed Hilbert space in the proof of [$MP_{2}$, Proposition 2], an odd analog of [$MP_{1}$, Proposition 1].
Elliptic operators on manifolds with boundary.   When the fibers $\mathcal{M}_{x}$ are compact manifolds with boundary, the situation becomes more complicated. In this case $\mathcal{D}_{x}$ is not even a closed operator usually. One needs to add an elliptic boundary condition $\mathcal{B}_{x}$ to $\mathcal{D}_{x}$, and then such a pair $(\mathcal{D}_{x},\mathcal{B}_{x})$ determines a regular operator with compact resolvents. Taking a local trivialization of $\mathcal{M}$ as above, one obtains a family of boundary value problems $(\mathcal{D}_{x},\mathcal{B}_{x})$ on a fixed manifold $M$. If both $\mathcal{D}_{x}$ and $\mathcal{B}_{x}$ depend continuously on $x$ in an appropriate sense, then the corresponding family of regular operators with compact resolvents is graph continuous. See [$P_{1}$, Appendix A.5] or [BZ, Corollary A.6.3]. Theorem C shows that such a family $\mathcal{D}$ can be made Riesz continuous by appropriate global trivializations of the corresponding Hilbert bundles.
Suppose now that we have a family of self-adjoint elliptic boundary value problems. As before, a local trivialization of $\mathcal{M}$ gives rise to a family of boundary value problems on a fixed manifold $M$. Unfortunately, there are no criteria of Riesz continuity for families of unbounded operators arising from boundary value problems, except for several special cases. Moreover, there are simple examples of such families that are not Riesz quasi-continuous, see e.g. Rellich’s example. Theorem E, however, requires only graph continuity of a family of unbounded operators, so it can be applied to families of elliptic self-adjoint boundary value problems, provided that they are neither essentially positive nor essentially negative and the space of parameters is good enough.
Acknowledgment.   I am grateful to N. V. Ivanov for attracting my attention to the work of Dixmier and Douady [DD] and for helpful remarks.
2 Preliminaries
Grassmanian.   The set $\operatorname{Gr}(H)$ of closed subspaces of $H$ is called the Grassmanian. An orthogonal projection may be identified with its range, which provides a bijection between $\operatorname{Gr}(H)$ and $\mathcal{P}(H)$. The Grassmanian is equipped with the topology (and even metric) induced by this bijection from the norm on the space $\mathcal{P}(H)$ of projections. We will use closed subspaces and projections by turn, depending on what is more convenient at the moment.
Canonical decomposition.   Recall a standard construction that will be used throughout the paper. The trivial Hilbert bundle over $\operatorname{Gr}(H)$ with the fiber $H$ is canonically decomposed into the direct sum
$\operatorname{Gr}(H)\times H=\mathcal{H}^{+}\oplus\mathcal{H}^{-}$
of two Hilbert bundles, whose fibers over $L\in\operatorname{Gr}$ are $\mathcal{H}^{+}_{L}=L$ and $\mathcal{H}^{-}_{L}=L^{\bot}$, the orthogonal complement of $L$ in $H$. This decomposition is locally trivial in the following sense: for every $L_{0}\in\operatorname{Gr}(H)$ there is a continuous map $g\colon B\to\mathcal{U}(H)$, where $B$ is the open ball of radius $1$ around $L_{0}$, such that $L=g_{L}L_{0}$ for every $L\in B$. In terms of projections, this is written as $p=g_{p}p_{0}g_{p}^{*}$ for $\left\|p-p_{0}\right\|<1$, where $p_{0}$ is the projection onto $L_{0}$. See the proof of [Ka, Theorem I.6.32] for the construction of such a map $g$.
An important subspace of $\operatorname{Gr}(H)$ is
$\operatorname{Gr}^{\star}(H)=\left\{L\in\operatorname{Gr}(H)\,|\,\dim L=\infty=\operatorname{codim}L\right\}.$
Let $L_{0}\in\operatorname{Gr}^{\star}(H)$ be a fixed subspace. The map $\mathcal{U}(H)\to\operatorname{Gr}^{\star}(H)$, $u\mapsto uL_{0}$ is a locally trivial principal bundle with the contractible structure group $\mathcal{U}(L_{0})\times\mathcal{U}(L_{0}^{\bot})$ over a paracompact base space and therefore has a section
(2.1) $g\colon\operatorname{Gr}^{\star}(H)\to\mathcal{U}(H),\;\;\text{so that}\;\;L=g_{L}L_{0}\;\;\text{for all}\;\;L\in\operatorname{Gr}^{\star}(H).$
It follows that the restrictions of $\mathcal{H}^{+}$ and $\mathcal{H}^{-}$ to $\operatorname{Gr}^{\star}$ are trivial Hilbert bundles with the structure group $\mathcal{U}$. In addition, contractibility of the total space $\mathcal{U}(H)$ and the fiber $\mathcal{U}\times\mathcal{U}$ implies contractibility of the base space $\operatorname{Gr}^{\star}$. See [AS, proof of Lemma 3.6].
Bounded transform.   Let us recall the definition and the main properties of the bounded transform. See, e.g., [Sc], Theorem 7.5 for the detail.
For a regular operator $A\colon H\to H^{\prime}$, the operator $1+A^{*}A$ is regular, self-adjoint, and surjective; its inverse $(1+A^{*}A)^{-1}$ is a positive bounded operator. The bounded transform (or the Riesz map)
$\bm{\chi}\colon\mathcal{R}(H,H^{\prime})\to\mathcal{B}(H,H^{\prime}),\quad\bm{\chi}(A)=A(1+A^{*}A)^{-1/2},$
defines the inclusion of the set $\mathcal{R}(H,H^{\prime})$ of regular operators to the closed unit ball
$\mathcal{D}(H,H^{\prime})=\left\{a\in\mathcal{B}(H,H^{\prime})\,|\,\left\|a\right\|\leqslant 1\right\}.$
The image of $\bm{\chi}$ consists of strict contractions, that is, operators $a\in\mathcal{D}$ such that $1-a^{*}a$ is injective; this image is dense in $\mathcal{D}$. For $a=\bm{\chi}(A)$ the following identities hold:
(2.2) $(1+A^{*}A)^{-1}=1-a^{*}a\;\;\text{and}\;\;A=a(1-a^{*}a)^{-1/2}.$
In particular, a regular operator $A$ has compact resolvents if and only if $\bm{\chi}(A)$ is essentially unitary. If a regular operator $A$ is self-adjoint, then so is $\bm{\chi}(A)$; more generally, $\bm{\chi}(A^{*})=\bm{\chi}(A)^{*}$.
Graphs of operators.   Let $\Gamma_{A}\subset H\oplus H^{\prime}$ denote the graph of a regular operator $A\colon H\to H^{\prime}$. There is a canonical isometric isomorphism
$\Gamma_{A}\to H,\quad\xi\oplus A\xi\mapsto\sqrt{1+A^{*}A}\,\xi\;\text{ for }\xi\in H.$
The inverse isomorphism $H\to\Gamma_{A}$, which we consider as an isometry $\mathbf{w}_{A}\colon H\to H\oplus H^{\prime}$ with the range $\Gamma_{A}$, is given by the formula
(2.3) $\mathbf{w}_{A}=(1+A^{*}A)^{-1/2}\oplus A(1+A^{*}A)^{-1/2}=(1+A^{*}A)^{-1/2}\oplus\bm{\chi}(A).$
The orthogonal projection onto $\Gamma_{A}$ is
(2.4) $\mathbf{p}_{A}=\mathbf{w}_{A}\mathbf{w}_{A}^{*}=\begin{pmatrix}(1+A^{*}A)^{-1}&(1+A^{*}A)^{-1}A^{*}\\ A(1+A^{*}A)^{-1}&1-(1+AA^{*})^{-1}\end{pmatrix}\in\mathcal{P}(H\oplus H^{\prime}).$
In terms of $a=\bm{\chi}(A)$ this is written as follows:
$\mathbf{w}_{A}=\tilde{\mathbf{w}}_{a}=\sqrt{1-a^{*}a}\oplus a,\quad\quad\mathbf{p}_{A}=\tilde{\mathbf{p}}_{a}=\tilde{\mathbf{w}}_{a}\tilde{\mathbf{w}}_{a}^{*}=\begin{pmatrix}1-a^{*}a&\sqrt{1-a^{*}a}\,a^{*}\\ a\sqrt{1-a^{*}a}&aa^{*}\end{pmatrix},$
so $\mathbf{p}$ factors through a continuous map $\tilde{\mathbf{p}}\colon\mathcal{D}(H,H^{\prime})\to\mathcal{P}(H\oplus H^{\prime})$. Restricting $\tilde{\mathbf{p}}$ to the subspace ${{}^{r}\mathcal{R}}$ of $\mathcal{D}$, we see that the identity map ${{}^{r}\mathcal{R}}\to{{}^{g}\mathcal{R}}$ is included into the following commutative square and thus is continuous:
[…]
Fredholm operators.   Similar to the bounded case, a regular operator is called Fredholm if its range is closed and its kernel and cokernel are finite-dimensional. The bounded transform $\bm{\chi}(A)$ is equal to the composition of the isomorphism $\mathbf{w}_{A}\colon H\to\Gamma_{A}$ and the restriction $p^{\prime}_{A}\colon\Gamma_{A}\to H^{\prime}$ of the orthogonal projection $p^{\prime}\colon H\oplus H^{\prime}\to H^{\prime}$. The range of $p^{\prime}_{A}$ coincides with the range of $A$. The kernel of $p^{\prime}_{A}$ is isomorphic to the kernel of $A$. Therefore, a regular operator $A$ is Fredholm if and only if $\bm{\chi}(A)$ is Fredholm.
Graph-to-strong* continuity of the bounded transform.   By the definition of Riesz topology, the bounded transform $\bm{\chi}\colon\mathcal{R}(H,H^{\prime})\to\mathcal{D}(H,H^{\prime})$ is Riesz-to-norm continuous. In the proof of Theorem A we also need another type of continuity of $\bm{\chi}$.
Recall that a family of bounded operators $a_{x}\in\mathcal{B}(H,H^{\prime})$ is said to be continuous in the strong operator topology (or strongly continuous) if the family $a_{x}\xi$ of vectors in $H^{\prime}$ is continuous for every $\xi\in H$. A family $a_{x}$ is said to be continuous in the strong* operator topology if both families $a_{x}$ and $a_{x}^{*}$ are strongly continuous.
Proposition 2.1.
The bounded transform is graph-to-strong* continuous.
Proof. Since $\mathbf{p}_{A^{*}}=1-J\mathbf{p}_{A}J^{*}$, where $J=\left(\begin{smallmatrix}0&-1\\ 1&0\end{smallmatrix}\right)$, the map $A\mapsto A^{*}$ is graph continuous. Since $\bm{\chi}(A)^{*}=\bm{\chi}(A^{*})$, it is enough to prove graph-to-strong continuity of $\bm{\chi}$.
For $A\in\mathcal{R}(H,H^{\prime})$ we denote
(2.5) $a=\bm{\chi}(A),\quad q=(1+A^{*}A)^{-1/2}=(1-a^{*}a)^{1/2},\quad b=aq.$
The projection $\mathbf{p}_{A}$ is given by the formula
(2.6) $\mathbf{p}_{A}=\tilde{\mathbf{p}}_{a}=\begin{pmatrix}1-a^{*}a&\sqrt{1-a^{*}a}\,a^{*}\\ a\sqrt{1-a^{*}a}&aa^{*}\end{pmatrix}=\begin{pmatrix}q^{2}&qa^{*}\\ aq&aa^{*}\end{pmatrix}=\begin{pmatrix}q^{2}&b^{*}\\ b&aa^{*}\end{pmatrix}.$
Since the map $A\mapsto b$ is given by the $(2,1)$-component of $\mathbf{p}_{A}$, it is graph-to-norm continuous. Similarly, the map $A\mapsto q$ is the square root of the $(1,1)$-component of $\mathbf{p}_{A}$ and thus is graph-to-norm continuous.
Choose $\xi\in H$ and $A_{0}\in\mathcal{R}(H,H^{\prime})$. Let $a_{0}$, $q_{0}$, $b_{0}$ be determined by $A_{0}$ as in (2.5). We need to show that the map ${{}^{g}\mathcal{R}}(H,H^{\prime})\to H^{\prime}$, $A\mapsto\bm{\chi}(A)\xi$ is continuous at $A=A_{0}$. Let $\varepsilon>0$. The operator $q_{0}$ is self-adjoint and injective and thus has dense range, so one can find $\eta\in H$ such that $\left\|q_{0}\eta-\xi\right\|<\varepsilon$. Let $V$ be a neighborhood of $A_{0}$ in ${{}^{g}\mathcal{R}}(H,H^{\prime})$ such that $\left\|(q-q_{0})\eta\right\|<\varepsilon$ and $\left\|(b-b_{0})\eta\right\|<\varepsilon$ for every $A\in V$. Then
$\left\|a\xi-a_{0}\xi\right\|\;<\;\left\|\left(a-a_{0}\right)q_{0}\eta\right\|+2\varepsilon\;\leqslant\;\left\|a\left(q_{0}-q\right)\eta\right\|+\left\|\left(aq-a_{0}q_{0}\right)\eta\right\|+2\varepsilon$
$\leqslant\;\left\|\left(q-q_{0}\right)\eta\right\|+\left\|\left(b-b_{0}\right)\eta\right\|+2\varepsilon<4\varepsilon$
for every $A\in V$. It follows that $\bm{\chi}$ is graph-to-strong continuous at an arbitrary point $A_{0}\in\mathcal{R}$ and thus on the whole $\mathcal{R}$. $\square$
Remark.
In fact, this reasoning provides a stronger result. It concerns another natural topology on $\mathcal{R}(H,H^{\prime})$, which is is strictly coarser then the graph topology, induced by the inclusion $\mathbf{p}\colon\mathcal{R}(H,H^{\prime})\hookrightarrow{{}^{s}{\mathcal{P}}}(H\oplus H^{\prime})$ from the strong operator topology on the space of projections. On the subset of self-adjoint operators it coincides with the strong resolvent topology, which is induced by the map $A\mapsto(A+i)^{-1}$, or equivalently by the map $A\mapsto(A-i)^{-1}$, from the strong operator topology on bounded operators. The reasoning in the proof of Proposition 2.1 shows that the bounded transform $\bm{\chi}\colon\mathcal{R}\to\mathcal{D}$ is continuous with respect to this coarser topology on $\mathcal{R}$ and the strong* topology on $\mathcal{D}$.
Proposition 2.1 follows from a similar result for self-adjoint operators and the restriction of $\bm{\chi}$ to $\mathcal{R}^{\mathrm{sa}}$, $A\mapsto A(1+A^{2})^{-1/2}$, by the standard trick of passing from $A$ to the self-adjoint operator $\left(\begin{smallmatrix}0&A^{*}\\ A&0\end{smallmatrix}\right)$. For self-adjoint operators, a similar reasoning provides continuity of an arbitrary function $f\in C_{b}(\mathbb{R})$ with respect to the strong resolvent topology on $\mathcal{R}^{\mathrm{sa}}$ and the strong topology on $\mathcal{D}$. A different proof of this fact based on the spectral theorem is given by [RS], Theorem VIII.20(b). A similar result for compact spaces of parameters is given by [W], Proposition 1.1 of Ch. Wahl.
Action of the unitary groups.   The unitary groups $\mathcal{U}(H)$ and $\mathcal{U}(H^{\prime})$ act on $\mathcal{R}(H,H^{\prime})$ by the right and left multiplication respectively. These actions preserve the subsets $\mathcal{R}_{F}$ and $\mathcal{R}_{K}$. For $u\in\mathcal{U}(H^{\prime})$ and $v\in\mathcal{U}(H)$ we have
$\bm{\chi}(uA)=u\bm{\chi}(A),\quad\bm{\chi}(Av)=\bm{\chi}(A)v,\quad\Gamma_{uA}=(1\oplus u)\Gamma_{A},\quad\Gamma_{Av}=(v^{-1}\oplus 1)\Gamma_{A},$
so the actions of $\mathcal{U}(H)$ and $\mathcal{U}(H^{\prime})$ (equipped with the norm topology) are continuous on both ${{}^{r}\mathcal{R}}$ and ${{}^{g}\mathcal{R}}$.
Recall that ${{}^{s}{\mathcal{U}}}(H)$, the unitary group equipped with the strong operator topology, is a topological group. Indeed, the multiplication map $\mathcal{B}(H)\times\mathcal{B}(H)\to\mathcal{B}(H)$, $(a,b)\mapsto ab$, is strongly continuous on $S\times\mathcal{B}(H)$ for every bounded subset $S$ of $\mathcal{B}(H)$. In particular, it is strongly continuous on $\mathcal{U}(H)\times\mathcal{U}(H)$. For every $u,u_{0}\in\mathcal{U}(H)$ and $\xi\in H$,
$\left\|(u^{-1}-u_{0}^{-1})\xi\right\|=\left\|u(u^{-1}-u_{0}^{-1})\xi\right\|=\left\|(u-u_{0})u_{0}^{-1}\xi\right\|,$
so the inverse map $\mathcal{U}(H)\to\mathcal{U}(H)$, $u\mapsto u^{-1}=u^{*}$, is also strongly continuous.
The actions of ${{}^{s}{\mathcal{U}}}(H)$ and ${{}^{s}{\mathcal{U}}}(H^{\prime})$ on ${{}^{r}\mathcal{R}}$ and ${{}^{g}\mathcal{R}}$ are not continuous. However, continuity of their actions on $\mathcal{K}(H\oplus H^{\prime})$ provides the following result.
Proposition 2.2.
The actions of the groups ${{}^{s}{\mathcal{U}}}(H)$ and ${{}^{s}{\mathcal{U}}}(H^{\prime})$ on ${{}^{g}\mathcal{R}}_{K}(H,H^{\prime})$ are continuous
Proof. The graph of an operator with compact resolvents belongs to the restricted Grassmanian
(2.7) $\operatorname{Gr}_{K}(H\oplus H^{\prime})\cong\mathcal{P}_{K}(H\oplus H^{\prime})=\left\{p\in\mathcal{P}(H\oplus H^{\prime})\,|\,p-p_{\infty}\text{ is compact}\right\},$
where $p_{\infty}\in\mathcal{P}(H\oplus H^{\prime})$ is the projection onto $0\oplus H^{\prime}$. The restriction of the embedding $\mathbf{p}\colon{{}^{g}\mathcal{R}}\hookrightarrow\mathcal{P}$ to operators with compact resolvents provides the embedding $\mathbf{p}\colon{{}^{g}\mathcal{R}}_{K}\hookrightarrow\mathcal{P}_{K}$. The left action of ${{}^{s}{\mathcal{U}}}(H^{\prime})$ on ${{}^{g}\mathcal{R}}(H,H^{\prime})$ lifts to the action on $\mathcal{P}(H\oplus H^{\prime})$:
$(u,p)\mapsto(1\oplus u)p(1\oplus u^{-1})=(1\oplus u)(p-p_{\infty})(1\oplus u)^{-1}+p_{\infty}.$
Since the action of ${{}^{s}{\mathcal{U}}}(H\oplus H^{\prime})$ on $\mathcal{K}(H\oplus H^{\prime})$ by conjugation is continuous, the action of ${{}^{s}{\mathcal{U}}}(H^{\prime})$ on $\mathcal{P}_{K}(H\oplus H^{\prime})$ is also continuous, as well as its restriction to ${{}^{g}\mathcal{R}}_{K}(H,H^{\prime})$. The right action of ${{}^{s}{\mathcal{U}}}(H)$ on ${{}^{g}\mathcal{R}}_{K}$ is continuous by the same argument. $\square$
3 Non-self-adjoint operators
[…]
we obtain a desired map $v\colon{{}^{g}\mathcal{R}}\to{{}^{s}{\mathcal{U}}}(H)$. If $\Gamma$ is considered as a Hilbert bundle over ${{}^{r}\mathcal{R}}$, then $\mathbf{w}$ becomes a trivialization in the category of bundles with the structure group $\mathcal{U}(H)$, so $v$ is continuous as a map ${{}^{r}\mathcal{R}}\to\mathcal{U}(H)$ as well.
Isometries.   Let
=$\mathcal{I}=\left\{u\in\mathcal{B}(H,H\oplus H^{\prime})\,|\,u^{*}u=\operatorname{Id}_{H}\right\}$
be the space of isometries (i.e., isometric embeddings) $H\to H\oplus H^{\prime}$ equipped with the usual norm topology. The canonical map
$\bm{\pi}\colon\mathcal{I}\to\operatorname{Gr}(H\oplus H^{\prime}),\quad u\mapsto\operatorname{Ran}(u).$
takes an isometry $u$ to its range; the corresponding projection is $uu^{*}$. The image $\bm{\pi}(\mathcal{I})$ of this map consists of infinite-dimensional subspaces of $H\oplus H^{\prime}$. We will need the following simple corollary of Kuiper’s theorem [Ku] about contractibility of $\mathcal{U}(H)$.
Lemma 3.1.
The map $\bm{\pi}\colon\mathcal{I}\to\bm{\pi}(\mathcal{I})$ is a principal $\mathcal{U}(H)$-bundle admitting a continuous section (and thus is trivial).
Proof. The group $\mathcal{U}(H)$ acts on $\mathcal{I}$ by the right multiplication. Every two isometries with the same range differ by an uniquely defined unitary operator $H\to H$, so this action turns $\bm{\pi}$ into a principal $\mathcal{U}(H)$-bundle. This bundle has a section by the same standard arguments as were used in the beginning of Section 2. Indeed, for every $u_{0}\in\mathcal{I}$ and $L_{0}=\bm{\pi}(u_{0})$ there is a continuous map $g$ from a neighborhood $B$ of $L_{0}$ to $\mathcal{U}(H\oplus H^{\prime})$ such that $L=g_{L}L_{0}$ for $L\in B$. The rule $L\mapsto g_{L}\cdot u_{0}$ determines a local section of $\bm{\pi}$ over $B$. Hence $\bm{\pi}$ is a locally trivial principal bundle over $\bm{\pi}(\mathcal{I})$ with the contractible structure group $\mathcal{U}(H)$. The base space $\bm{\pi}(\mathcal{I})$ is metric and thus paracompact. Therefore, $\bm{\pi}$ has a global section. $\square$
Proof of Theorem A.   The graphs of regular operators have infinite dimension, so ${{}^{g}\mathcal{R}}\subset\bm{\pi}(\mathcal{I})\subset\operatorname{Gr}(H\oplus H^{\prime})$. By Lemma 3.1, the restriction of $\bm{\pi}$ to ${{}^{g}\mathcal{R}}$ has a section. For the rest of the proof we fix such a section $u\colon{{}^{g}\mathcal{R}}\to\mathcal{I}$. In other words, $u$ is a graph-to-norm continuous map taking a regular operator $A$ to an isometry $u_{A}\colon H\to H\oplus H^{\prime}$ with the range $\Gamma_{A}$.
Let $p^{\prime}\colon H\oplus H^{\prime}\to H^{\prime}$ be the projection onto the second summand. The composition
$\varphi_{A}=p^{\prime}\cdot u_{A}\colon H\to H^{\prime}$
is a contraction, and the corresponding map $\varphi\colon{{}^{g}\mathcal{R}}(H,H^{\prime})\to\mathcal{D}(H,H^{\prime})$ is continuous.
The maps defined above are not canonical: they depend on the choice of a section $u$. On the other hand, formula (2.3) provides a canonical isometry $\mathbf{w}_{A}\colon H\to H\oplus H^{\prime}$ with the range $\Gamma_{A}$:
$\mathbf{w}_{A}=(1+A^{*}A)^{-1/2}\oplus\bm{\chi}(A)=\sqrt{1-a^{*}a}\oplus a,\;\;\text{where}\;\;a=\bm{\chi}(A).$
The corresponding contraction $p^{\prime}\cdot\mathbf{w}_{A}\colon H\to H^{\prime}$ coincides with the bounded transform $\bm{\chi}(A)$ of $A$. In contrast with $\varphi$, it is not continuous as a map ${{}^{g}\mathcal{R}}\to\mathcal{D}$.
Two isometries $u_{A},\mathbf{w}_{A}\colon H\to H\oplus H^{\prime}$ have the same range $\Gamma_{A}$ and thus differ by a unitary operator
$v_{A}=\mathbf{w}_{A}^{*}\cdot u_{A}\in\mathcal{U}(H),\;\;\text{so that}\;\;u_{A}=\mathbf{w}_{A}\cdot v_{A}.$
We claim that the map $v\colon\mathcal{R}(H,H^{\prime})\to\mathcal{U}(H)$ defined by this formula satisfies conclusions of the theorem. Indeed, for $\Phi_{A}=A\cdot v_{A}$ we have
$\bm{\chi}(\Phi_{A})=\bm{\chi}(A\cdot v_{A})=\bm{\chi}(A)\cdot v_{A}=p^{\prime}\cdot\mathbf{w}_{A}\cdot v_{A}=p^{\prime}\cdot u_{A}=\varphi_{A},$
so $\bm{\chi}\circ\Phi=\varphi\colon\mathcal{R}(H,H^{\prime})\to\mathcal{B}(H,H^{\prime})$ is graph-to-norm continuous and thus $\Phi$ is graph-to-Riesz continuous. By the construction, $u\colon\mathcal{R}\to\mathcal{I}$ is graph-to-norm (and thus also Riesz-to-norm) continuous. The coisometry $\mathbf{w}_{A}^{*}\colon H\oplus H^{\prime}\to H$ is given by the formula
$\mathbf{w}_{A}^{*}(\xi\oplus\eta)\;=\;(1+A^{*}A)^{-1/2}\;\xi+\bm{\chi}(A)^{*}\,\eta\;=\;\sqrt{1-a^{*}a}\;\xi+a^{*}\,\eta\;\;\;\text{for}\;\;\xi\in H,\;\eta\in H^{\prime}.$
Clearly, the maps $A\mapsto\mathbf{w}_{A}^{*}$ and $A\mapsto v_{A}=\mathbf{w}_{A}^{*}\cdot u_{A}$ are Riesz-to-norm continuous.
The first component of $\mathbf{w}^{*}$, $A\mapsto(1+A^{*}A)^{-1/2}$, is the square root of the $(1,1)$-component of $\mathbf{p}_{A}$ and thus is even graph-to-norm continuous. The second component $A\mapsto\bm{\chi}(A)^{*}$ is graph-to-strong continuous by Proposition 2.1. Therefore, the map $A\mapsto\mathbf{w}_{A}^{*}$ is graph-to-strong continuous. The multiplication map $(b,c)\mapsto b\cdot c$ is strongly continuous on uniformly bounded subsets of operators, and the norms of $\mathbf{w}_{A}^{*}$ and $u_{A}$ are bounded by $1$. It follows that the map $A\mapsto v_{A}=\mathbf{w}_{A}^{*}\cdot u_{A}$ is graph-to-strong continuous. This completes the proof of the theorem. $\square$
Remark.
Our construction of $v$ depends on the choice of a section $u$. Since $\bm{\pi}$ is a principal $\mathcal{U}(H)$-bundle, any other choice of a section differs from $u$ by the multiplication by a graph-to-norm continuous map ${{}^{g}\mathcal{R}}(H,H^{\prime})\to\mathcal{U}(H)$. Therefore, our construction determines $v$ up to the right multiplication by such a map.
Proof of Theorem B.   In order to prove the first part of the theorem we need to show that $\Phi$ is homotopic to the identity as a self-map of both ${{}^{r}\mathcal{R}}_{K}$ and ${{}^{g}\mathcal{R}}_{K}$. For the second part we need to prove a similar result for ${{}^{r}\mathcal{R}}_{F}$ and ${{}^{g}\mathcal{R}}_{F}$.
By Kuiper’s theorem, the group $\mathcal{U}(H)$ is contractible in the norm topology, so the map $v\colon\mathcal{U}\to\mathcal{U}$ is homotopic to the identity map $\mathcal{U}\to\mathcal{U}$. Since the action of $\mathcal{U}(H)$ on ${{}^{r}\mathcal{R}}(H,H^{\prime})$ is continuous, a homotopy between these two maps provides a homotopy between $\Phi\colon{{}^{r}\mathcal{R}}_{K}\to{{}^{r}\mathcal{R}}_{K}$ and $\operatorname{Id}\colon{{}^{r}\mathcal{R}}_{K}\to{{}^{r}\mathcal{R}}_{K}$, and a similar homotopy for ${{}^{r}\mathcal{R}}_{F}\to{{}^{r}\mathcal{R}}_{F}$.
By a result of Dixmier and Douady [DD, Lemma 3], the group ${{}^{s}{\mathcal{U}}}(H)$ is contractible, so the map $v\colon{{}^{s}{\mathcal{U}}}\to{{}^{s}{\mathcal{U}}}$ is homotopic to the identity map ${{}^{s}{\mathcal{U}}}\to{{}^{s}{\mathcal{U}}}$. The action of ${{}^{s}{\mathcal{U}}}(H)$ on ${{}^{g}\mathcal{R}}_{K}(H,H^{\prime})$ is continuous (see Proposition 2.2), so a homotopy between these two maps provides a homotopy between the maps $\Phi\colon{{}^{g}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$ and $\operatorname{Id}\colon{{}^{g}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$. Therefore, the restriction of $\Phi$ to the subspace of operators with compact resolvents is homotopy inverse to the identity map ${{}^{r}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$.
In order to finish the proof of the Fredholm part, we need to show that the maps $\Phi\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$ and $\operatorname{Id}\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$ are homotopic. This part is more delicate because the action of ${{}^{s}{\mathcal{U}}}$ on ${{}^{g}\mathcal{R}}_{F}$ is not continuous, so we cannot directly apply a contraction of ${{}^{s}{\mathcal{U}}}$ for this purpose. However, we can use a part of [$P_{3}$, Theorem A] which states that the embedding ${{}^{g}\mathcal{R}}_{K}\hookrightarrow{{}^{g}\mathcal{R}}_{F}$ is a homotopy equivalence. We already proved that the maps $\Phi\colon{{}^{g}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$ and $\operatorname{Id}\colon{{}^{g}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$ are homotopic; it follows that $\Phi\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$ and $\operatorname{Id}\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$ are also homotopic. Therefore, the restriction of $\Phi$ to the subspace of Fredholm operators is homotopy inverse to the identity map ${{}^{r}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$. $\square$
Proof of Theorem C.   Suppose that a global trivialization of $\mathcal{H}^{\prime}$ is fixed. Choose an arbitrary trivialization $\tau\colon\mathcal{H}\to X\times H$ of $\mathcal{H}$. In these trivializations $\mathcal{A}$ becomes a graph continuous map $X\to\mathcal{R}_{K}(H,H^{\prime})$. Let $v\colon{{}^{g}\mathcal{R}}(H,H^{\prime})\to{{}^{s}{\mathcal{U}}}(H)$ be a map provided by Theorem A. Then the map $X\to{{}^{s}{\mathcal{U}}}(H)$, $x\mapsto v(\mathcal{A}_{x})$ is continuous, so
$\tau_{v}\colon(x,\xi)\mapsto(x,v(\mathcal{A}_{x})\xi)$
is an automorphism of the trivial Hilbert bundle $X\times H$. The inverse automorphism $\tau_{v}^{-1}$ takes $\mathcal{A}$ to a Riesz continuous map
$\mathcal{A}^{\prime}\colon X\to\mathcal{R}_{K}(H,H^{\prime}),\quad\mathcal{A}^{\prime}_{x}=\mathcal{A}_{x}\cdot v(\mathcal{A}_{x}).$
It follows that the composition $\tau_{v}^{-1}\circ\tau$ satisfies conclusion of the theorem.
If instead a global trivialization of $\mathcal{H}$ is fixed, then one can interchange $\mathcal{H}$ and $\mathcal{H}^{\prime}$ by passing to the adjoint operators. Since the map $A\mapsto A^{*}$ is both graph and Riesz continuous, it follows from the first part of the proof that there is a global trivialization of $\mathcal{H}^{\prime}$ taking $\mathcal{A}$ to a Riesz continuous map. $\square$
$K^{0}$ index for graph continuous families.   By a classical result of Atiyah and Jänich, the space $\mathcal{B}_{F}(H)$ of bounded Fredholm operators is a classifying space for the functor $K^{0}$. See [A], Theorem A1. For an arbitrary parameter space $X$, $K^{0}(X)$ might be defined as the set $[X,\mathcal{B}_{F}]$ of homotopy classes of maps $X\to\mathcal{B}_{F}(H)$, with the group structure on $[X,\mathcal{B}_{F}]$ introduced in the standard way. The class of a continuous map $\alpha\colon X\to\mathcal{B}_{F}$ in $K^{0}(X)$ is called the index of $\alpha$:
$\operatorname{ind}(\alpha)=[\alpha]\in[X,\mathcal{B}_{F}]=K^{0}(X).$
It can be easily seen, using linear homotopies, that the bounded transform map $\bm{\chi}\colon{{}^{r}\mathcal{R}}_{F}\to\mathcal{B}_{F}$ is homotopy inverse to the embedding $\mathcal{B}_{F}\hookrightarrow{{}^{r}\mathcal{R}}_{F}$, so both these maps are homotopy equivalencies. Therefore, the index of a Riesz continuous family of Fredholm operators is naturally defined as the index of its bounded transform:
$\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\bm{\chi}\circ\mathcal{A})\in[X,\mathcal{B}_{F}]=K^{0}(X)\;\;\text{for}\;\;\mathcal{A}\colon X\to{{}^{r}\mathcal{R}}_{F}.$
Let now $\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}_{F}$ be a graph continuous family of Fredholm operators. The space ${{}^{g}\mathcal{R}}_{F}$ is also a classifying space for the functor $K^{0}$, see [J], Theorem 3.5(i) or [$P_{3}$], Theorem A. Hence the homotopy class of $\mathcal{A}$ determines an element of $K^{0}(X)$ which can be called the index of $\mathcal{A}$:
(3.1) $\operatorname{ind}(\mathcal{A})=[\mathcal{A}]\in[X,{{}^{g}\mathcal{R}}_{F}]=K^{0}(X).$
By the definition, two graph continuous families of Fredholm operators are homotopic if and only if their indices coincide.
Theorems A and B of the present paper allow to relate the index of $\mathcal{A}$ given by (3.1) to the Fredholm index of bounded operators as follows. The composition of $\mathcal{A}$ with a map $\Phi\colon{{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ provided by Theorem A is Riesz continuous, so $\bm{\chi}\circ\Phi\circ\mathcal{A}$ is a norm continuous family of bounded Fredholm operators. Theorem B provides the equality
$\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\Phi\circ\mathcal{A})=\operatorname{ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})\in[X,\mathcal{B}_{F}]=K^{0}(X)\;\;\text{for}\;\;\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}_{F},$
which can be used as the definition of the index of $\mathcal{A}$. Note that Theorem B implies that the homotopy class of $\Phi\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{r}\mathcal{R}}_{F}$ does not depend on the choice of $v$, so $\operatorname{ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})$ is also independent of this choice. Equivalently, the index can be defined in terms of graphs of operators as described in the Introduction.
For a Riesz continuous family $\mathcal{A}\colon X\to\mathcal{R}_{F}$ (in particular, for a norm continuous family of bounded Fredholm operators) the two definitions of the index coincide. Indeed, $\Phi(\mathcal{A}_{x})=\mathcal{A}_{x}\cdot v(\mathcal{A}_{x})$, so $\Phi\circ\mathcal{A}$ differs from $\mathcal{A}$ by the multiplication by a map $v\circ\mathcal{A}\colon X\to\mathcal{U}(H)$. If $\mathcal{A}$ is Riesz continuous, then $v\circ\mathcal{A}$ is norm continuous. Since $\mathcal{U}(H)$ is contractible, $\Phi\circ\mathcal{A}$ and $\mathcal{A}$ are homotopic as maps from $X$ to ${{}^{r}\mathcal{R}}_{F}$ and thus their indices coincide.
Recently N. Ivanov [$I_{1}$, $I_{2}$] developed a new approach to the analytical index of families. He introduced the notion of a “Fredholm family” and defined the $K^{0}(X)$-valued index, which we will denote by $\operatorname{Ind}$, for Fredholm families parametrized by points of a paracompact space $X$. See [$I_{2}$], Section 3. A Fredholm family is a family of Fredholm operators $\mathcal{A}_{x}$ such that the finite-dimensional subspaces $\operatorname{\mathbb{1}}_{[0,\varepsilon]}(|\mathcal{A}_{x}|)$ and $\operatorname{\mathbb{1}}_{[0,\varepsilon]}(|\mathcal{A}^{*}_{x}|)$, as well as the restriction of $\mathcal{A}_{x}$ to $\operatorname{\mathbb{1}}_{[0,\varepsilon]}(|\mathcal{A}_{x}|)$, depend continuously on $x$ for an appropriate locally chosen $\varepsilon>0$, without any additional continuity assumptions. Ivanov’s index $\operatorname{Ind}$ is defined both for operators acting on a fixed Hilbert space and on fibers of a Hilbert bundle; it is invariant under isomorphisms of Hilbert bundles and under passing to the bounded transform.
Let $\mathcal{A}$ be a graph continuous family of Fredholm operators in $H$ parametrized by points of a paracompact space $X$. Then $\mathcal{A}$ is a Fredholm family in Ivanov’s sense, and we claim that our definition of the index of $\mathcal{A}$ provides the same element of $K^{0}(X)$ as Ivanov’s definition. Indeed, the properties of $\operatorname{Ind}$ mentioned above imply the equality $\operatorname{Ind}(\mathcal{A})=\operatorname{Ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})$. Consider the norm continuous family $\alpha=\bm{\chi}\circ\Phi\circ\mathcal{A}$ of bounded Fredholm operators in $H$. By Theorem 7.2 of [$I_{2}$], $\operatorname{Ind}\alpha$ coincides with the index of $\alpha$ defined by Segal in [Se]. The last index is equal to the homotopy class $[\alpha]$ of $\alpha$ in $[X,\mathcal{B}_{F}(H)]=K^{0}(X)$ by [Se], Theorem 5.1(b). Taking all this together, we obtain
$\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\alpha)=[\alpha]=\operatorname{Ind}(\alpha)=\operatorname{Ind}(\mathcal{A})\in K^{0}(X).$
4 Polarizations
Polarizations of a Hilbert space.   A decomposition of a Hilbert space $H$ into the orthogonal sum $H=H^{+}\oplus H^{-}$ of two closed infinite-dimensional subspaces is called a polarization of $H$. It is determined by the subspace $H^{+}$ or equivalently by the orthogonal projection $p\in\mathcal{P}^{\star}(H)$ onto $H^{+}$; we will identify polarizations with the corresponding subspaces/projections.
There are two natural equivalence relations for polarizations of $H$, which we call “finite” and “compact”. Two polarizations $p$ and $q$ of $H$ are said to be compactly compatible if $p-q$ is a compact operator; they are said to be finitely compatible if $p-q$ is a finite rank operator (that is, the corresponding subspaces are commensurable).
The compact restricted Grassmanian corresponding to the polarization given by a projection $p$ is defined as the space of all polarizations compatible with $p$:
$\operatorname{Gr}_{K}=\operatorname{Gr}_{K}(H;p)=\left\{q\in\mathcal{P}^{\star}(H)\,|\,q-p\in\mathcal{K}(H)\right\}.$
It is equipped with the norm topology induced from the natural inclusion $\operatorname{Gr}_{K}(H;p)\hookrightarrow\mathcal{P}(H)$, or equivalently the inclusion $\operatorname{Gr}_{K}(H;p)\hookrightarrow\mathcal{K}(H)$, $q\mapsto q-p$. The finite restricted Grassmanian $\operatorname{Gr}_{f}(H;p)$ is defined in a similar manner; it is a subspace of $\operatorname{Gr}_{K}(H;p)$. The restricted Grassmanian depends only on the corresponding equivalence class of the projection $p$.
Subbundles of a Hilbert bundle.   We will consider only locally trivial Hilbert bundles with a separable fiber $H$. However, we want to deal with their subbundles which are not necessarily locally trivial. Such subbundles arise, for example, as spectral sections of graph continuous families of Fredholm operators.
In general, a subbundle $\mathcal{E}$ of $\mathcal{H}$ is given by a family of subspaces $\mathcal{E}_{x}\subset\mathcal{H}_{x}$ which is continuous in an appropriate sense. We will use the following notion of continuity: we say that $\mathcal{E}$ is a subbundle of $\mathcal{H}$ if $\mathcal{E}$ is the range of a bundle projection, that is, a continuous bundle map $p\colon\mathcal{H}\to\mathcal{H}$ such that $p_{x}\in\mathcal{P}(\mathcal{H}_{x})$ for every $x\in X$. In this case, $\mathcal{H}\ominus\mathcal{E}=\operatorname{Ran}(1-p)$ is also a subbundle. If both $\mathcal{E}$ and $\mathcal{H}\ominus\mathcal{E}$ are subbundles of $\mathcal{H}$ in the sense of [DD], then they are subbundles in our sense. See [DD], Proposition 16.
Equivalently, for every $x\in X$, some (and then every) local trivialization of $\mathcal{H}$ over a neighborhood $X^{\prime}$ of $x$ takes $p$ to a continuous map $X^{\prime}\to{{}^{s}{\mathcal{P}}}(H)$, where ${{}^{s}{\mathcal{P}}}(H)$ denotes the space $\mathcal{P}(H)$ equipped with the strong operator topology. In particular, subbundles of the trivial Hilbert bundle $H_{X}=X\times H$ are in one-to-one correspondence with continuous maps $X\to{{}^{s}{\mathcal{P}}}(H)$.
Two subbundles of $\mathcal{H}$ given by bundle projections $p$ and $q$ are said to be (compactly) compatible if $p-q$ is a norm continuous family of compact operators. They are said to be finitely compatible if additionally all operators $p_{x}-q_{x}$ have finite rank.
Polarizations of a Hilbert bundle.   We need only compact polarizations in this paper, so we will mostly use the term “polarization” without the adjective “compact”. The changes required to obtain the finite version are obvious.
Locally, a polarization of $\mathcal{H}$ is determined by a decomposition of $\mathcal{H}$, over some open subset $X^{\prime}\subset X$, into the orthogonal sum $\mathcal{H}=\mathcal{H}^{+}\oplus\mathcal{H}^{-}$ of two subbundles of infinite rank. Equivalently, it is determined by the subbundle $\mathcal{H}^{+}$ of $\left.\mathcal{H}\right|_{X^{\prime}}$, or by the bundle projection with the range $\mathcal{H}^{+}$. By definition, two compatible subbundles of $\mathcal{H}$ over $X^{\prime}$ determine the same polarization.
Globally, a polarization of $\mathcal{H}$ is determined by an open covering $(X_{i})$ of $X$ and subbundles $\mathcal{H}^{+}_{i}$ of $\left.\mathcal{H}\right|_{X_{i}}$ such that $\mathcal{H}^{+}_{i}$ has infinite rank and corank and is compatible with $\mathcal{H}^{+}_{j}$ over $X_{i}\cap X_{j}$ for every $i,j$. Following Ivanov [$I_{2}$], we call such a collection $(X_{i},\mathcal{H}^{+}_{i})$ a polarization atlas. Two polarization atlases are said to be equivalent if their union is also a polarization atlas. A polarization of $\mathcal{H}$ is an equivalence class of polarization atlases.
A polarization $\Pi$ of $\mathcal{H}$ determines the restricted Grassmanian bundle $\operatorname{Gr}_{K}(\mathcal{H};\Pi)$ in an obvious way. This is one of the motivations behind the notion of a polarization.
Two polarizations of $\mathcal{H}$ are said to be homotopic if they are restrictions to $X\times\left\{0\right\}$ and $X\times\left\{1\right\}$ of some polarization of the Hilbert bundle ×$\mathcal{H}\times[0,1]\to X\times[0,1]$.
Adapted trivializations.   Following [$I_{2}$], we say that a trivialization $\mathcal{H}\to X\times H$ is adapted to a subbundle $\mathcal{E}$ of $\mathcal{H}$ if it takes the family of subspaces $\mathcal{E}_{x}\subset\mathcal{H}_{x}$ to a continuous map $X\to\operatorname{Gr}(H)$. If a trivialization of $\mathcal{H}$ is adapted to $\mathcal{E}$, then it is adapted to every subbundle compatible with $\mathcal{E}$.
Let $\Pi$ be a polarization given by an atlas $(X_{i},\mathcal{E}_{i})$. We say that a trivialization of $\mathcal{H}$ is adapted to $\Pi$ if its restrictions to $X_{i}$ are adapted to $\mathcal{E}_{i}$. The notion of an adapted local trivialization is defined in an obvious way. We call a polarization $\Pi$ of $\mathcal{H}$ tame if it admits local adapted trivializations (that is, for every point $x\in X$ there is a local trivialization of $\mathcal{H}$ adapted to $\Pi$ over some neighborhood of $x$).
The notion of a polarization introduced by Ivanov in the appendix to the third version of his preprint [$I_{2}$] is a particular case of a tame finite polarization in our terminology.
Theorem 4.1.
Let $\Pi$ be a tame polarization of a locally trivial Hilbert bundle $\mathcal{H}$ over a paracompact space $X$. Then $\Pi$ admits a global adapted trivialization. Moreover, every adapted trivialization of $\mathcal{H}$ over a closed subset $Y$ of $X$ can be extended to a global adapted trivialization. In particular, every two adapted trivializations are homotopic (as trivializations adapted to $\Pi$).
Proof. Let us describe the plan of the proof. For an arbitrary polarization $\Pi$ we define a fiber bundle $\mathcal{T}=\mathcal{T}_{\Pi}$ over $X$ whose (local or global) sections are exactly (local or global) trivializations of $\mathcal{H}$ adapted to $\Pi$. In particular, $\mathcal{T}$ admits local sections if and only if $\Pi$ is tame. We will show that in this case the bundle $\mathcal{T}$ is locally trivial and has a contractible fiber. Such a bundle, over a paracompact base space $X$, admits global sections, and every section over a closed subset $Y\subset X$ can be extended to a global section [DD, Lemma 4]. Therefore, $\Pi$ admits global adapted trivializations, and every trivialization of $\mathcal{H}$ over $Y$ adapted to $\Pi$ can be extended to a global adapted trivialization.
Let $H_{X}$ be the trivial Hilbert bundle over $X$ with the fiber $H$. Consider the fiber bundle $\mathcal{U}(\mathcal{H},H_{X})$ associated with $\mathcal{H}$ whose sections are Hilbert bundle isomorphisms $\mathcal{H}\to H_{X}$ (that is, trivializations of $\mathcal{H})$. Every two such isomorphisms differ by an automorphism of $H_{X}$, which is defined by a continuous map $X\to{{}^{s}{\mathcal{U}}}(H)$. Therefore, $\mathcal{U}(\mathcal{H},H_{X})$ is a locally trivial principal bundle over $X$ with the structure group ${{}^{s}{\mathcal{U}}}(H)$. Its fiber over $x$ is ${{}^{s}{\mathcal{U}}}(\mathcal{H}_{x},H)$, and ${{}^{s}{\mathcal{U}}}(H)$ acts on the fiber by the left multiplication. A trivialization $\tau\colon\mathcal{H}\to H_{X}$ takes $\mathcal{U}(\mathcal{H},H_{X})$ to the trivial bundle $\mathcal{U}(H_{X},H_{X})=X\times{{}^{s}{\mathcal{U}}}(H)$.
The key ingredient of our proof is the bundle $\mathcal{T}\to X$ which we will now define. The total space $\mathcal{T}$ coincides with $\mathcal{U}(\mathcal{H},H_{X})$ as a set, and the projection to $X$ is the same, but we equip $\mathcal{T}$ with a finer topology determined by $\Pi$. Hence the identity map
(4.1) $\iota\colon\mathcal{T}\to\mathcal{U}(\mathcal{H},H_{X})$
will be continuous and, in particular, the projection $\mathcal{T}\to X$ will be continuous.
Suppose first that the polarization $\Pi$ is defined by a subbundle $\mathcal{E}$ of $\mathcal{H}$, that is, the atlas of $\Pi$ has only one chart $(X,\mathcal{E})$. Then $\mathcal{E}$ determines the map
(4.2) $\iota_{\mathcal{E}}\colon\mathcal{T}\to\operatorname{Gr}(H)\cong\mathcal{P}(H),\quad(x,u)\mapsto u(\mathcal{E}_{x})\;\;\text{for}\;\;x\in X\;\;\text{and}\;\;u\in\mathcal{U}(\mathcal{H}_{x},H).$
We equip $\mathcal{T}=\mathcal{T}_{\mathcal{E}}$ with the weakest topology for which both maps (4.1) and (4.2) are continuous.
If $\mathcal{E}$ is replaced by a compatible subbundle $\mathcal{F}$, then $\mathcal{T}_{\mathcal{F}}=\mathcal{T}_{\mathcal{E}}$. Indeed, in this case the difference $k_{x}=p_{x}-q_{x}$ of projections onto $\mathcal{F}_{x}$ and $\mathcal{E}_{x}$ is a norm continuous family of compact operators, so $\iota_{\mathcal{F}}$ differs from $\iota_{\mathcal{E}}$ by a continuous map
$\mathcal{U}(\mathcal{H},H_{X})\to\mathcal{K}(H),\quad(x,u)\mapsto uk_{x}u^{*}.$
Therefore, $\iota_{\mathcal{F}}$ and $\iota_{\mathcal{E}}$ induce the same refinement of the original topology on $\mathcal{U}(\mathcal{H},H_{X})$ and thus $\mathcal{T}_{\mathcal{E}}$ depends only on the equivalence class of compatible subbundles $\mathcal{E}$. This allows to define the bundle $\mathcal{T}_{\Pi}$ for an arbitrary polarization $\Pi$ of $\mathcal{H}$.
More precisely, let $\Pi$ be given by an atlas $(X_{i},\mathcal{E}_{i})$, $i\in I$. Then the restrictions of $\mathcal{T}_{\mathcal{E}_{i}}$ and $\mathcal{T}_{\mathcal{E}_{j}}$ to $X_{i}\cap X_{j}$ coincide. We equip $\mathcal{T}_{\Pi}$ with the topology defined by the requirement that its restriction to $X_{i}$ coincides with $\mathcal{T}_{\mathcal{E}_{i}}$ for every $i\in I$.
By the definition of $\mathcal{T}_{\Pi}$, its (local) sections are exactly (local) trivializations of $\mathcal{H}$ adapted to $\Pi$. In particular, $\mathcal{T}_{\Pi}$ admits local sections if $\Pi$ is tame. The following lemma provides a stronger result.
Lemma 4.2.
If $\Pi$ is tame, then the bundle $\mathcal{T}_{\Pi}$ is locally trivial.
Proof. Since $\Pi$ is tame, each point $x_{0}\in X$ has a neighborhood $X^{\prime}$ such that some trivialization of $\mathcal{H}$ over $X^{\prime}$ is adapted to $\Pi$. We can suppose that $X^{\prime}$ is contained in some $X_{i}$; then this trivialization takes $\mathcal{E}_{i}$ to the subbundle $\mathcal{E}^{\prime}$ given by a continuous map $\mathcal{E}^{\prime}\colon X^{\prime}\to\operatorname{Gr}^{\star}(H)$. Let $g\colon\operatorname{Gr}^{\star}\to\mathcal{U}(H)$ be a map given by (2.1) and $\bar{g}_{L}=g_{L}^{-1}$. Then the composition of $\mathcal{E}^{\prime}$ with $\bar{g}\colon\operatorname{Gr}^{\star}\to\mathcal{U}(H)$ provides the automorphism $G$ of the trivial Hilbert bundle $H_{X^{\prime}}$ which takes $\mathcal{E}^{\prime}_{x}$ to the subspace $E$ of $H$ independent of $x$. We can then replace the original adapted trivialization by its composition with $G$. The new trivialization takes $\mathcal{E}_{i}$ to the subbundle $E_{X^{\prime}}$ of $H_{X^{\prime}}$ and thus provides a bundle isomorphism $\left.\mathcal{T}\right|_{X^{\prime}}\to\mathcal{T}^{\prime}$, where the bundle $\mathcal{T}^{\prime}$ is build (in the way described above) from the trivial bundle $H_{X^{\prime}}$ and its subbundle $E_{X^{\prime}}$. We will prove that the bundle $\mathcal{T}^{\prime}$ is trivial and find its fiber.
By the definition, the total space $\mathcal{T}^{\prime}$ is obtained from the product $X^{\prime}\times{{}^{s}{\mathcal{U}}}(H)$ by refining its topology so that the map $\mathcal{T}^{\prime}\to\operatorname{Gr}^{\star}(H)$, $(x,u)\mapsto uE$ becomes continuous. This map does not actually depend on $x$. Therefore, $\mathcal{T}^{\prime}$ coincides with the trivial bundle $X^{\prime}\times T\to X^{\prime}$, where the fiber $T$ is obtained from ${{}^{s}{\mathcal{U}}}(H)$ by refining its topology so that
(4.3) $\iota_{E}\colon T\to\operatorname{Gr}^{\star}(H),\quad u\mapsto uE$
becomes continuous. In other words, $T$ has the topology induced by the inclusion
$T\hookrightarrow{{}^{s}{\mathcal{U}}}(H)\times\operatorname{Gr}^{\star}(H),\quad u\mapsto(u,uE).$
Hence the restriction of $\mathcal{T}_{\Pi}$ to a neighborhood of an arbitrary point $x_{0}$ is a trivial bundle with the fiber $T$. $\square$
Lemma 4.3.
The space $T$ defined above is contractible.
Proof. The fiber of $\iota_{E}$ over $E$ is the subspace of ${{}^{s}{\mathcal{U}}}(H)$ consisting of unitary operators preserving $E$. Denote this subspace by ${{}^{s}{\mathcal{U}}}_{E}$. We will show that (4.3) is a trivial bundle over $\operatorname{Gr}^{\star}$ and construct its trivialization $\varphi\colon{{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}\to T$.
Let $g\colon\operatorname{Gr}^{\star}(H)\to\mathcal{U}(H)$ be a continuous map such that $L=g_{L}E$ for all $L\in\operatorname{Gr}^{\star}(H)$, as in (2.1). We define $\varphi$ by the formula
$\varphi\colon{{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}(H)\;\to\;T,\quad\varphi(v,L)=g_{L}\cdot v.$
Since
(4.4) $(g_{L}v)E=g_{L}E=L\;\;\text{for every}\;\;v\in{{}^{s}{\mathcal{U}}}_{E},$
the composition $\iota_{E}\circ\varphi\colon{{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}\to\operatorname{Gr}^{\star}$ is the projection of the product on the second factor; in particular, it is continuous. The composition of $\varphi$ with the identity map $T\to{{}^{s}{\mathcal{U}}}$ is given by $(v,L)\mapsto g_{L}v$ and is also continuous. Therefore, $\varphi$ itself is continuous.
Equality (4.4) allows to recover $L$ and $v$ from $u=g_{L}v$ by taking $L=uE$ and $v=g_{L}^{-1}u$. Since $vE=g_{L}^{-1}uE=g_{L}^{-1}L=E$, the unitary operator $v=g_{L}^{-1}u$ belongs to ${{}^{s}{\mathcal{U}}}_{E}$, and we obtain the continuous map
$\psi\colon T\to{{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}(H),\quad\psi(u)=\left(g_{uE}^{-1}\cdot u,\,uE\right).$
By the construction, $\psi\circ\varphi(v,L)=(v,L)$ for every $(v,L)\in{{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}$. Conversely,
$\varphi\circ\psi(u)=\varphi\left(g_{uE}^{-1}\cdot u,\,uE\right)=g_{uE}\cdot g_{uE}^{-1}\cdot u=u\;\;\text{for every}\;\;u\in T.$
Therefore, $\psi$ is inverse to $\varphi$ and thus $\varphi$ is a homeomorphism.
The domain of $\varphi$ is the product of two factors. The second factor $\operatorname{Gr}^{\star}$ is contractible, see [AS, proof of Lemma 3.6]. The first factor ${{}^{s}{\mathcal{U}}}_{E}={{}^{s}{\mathcal{U}}}(E)\times{{}^{s}{\mathcal{U}}}(E^{\bot})$ is the product of two copies of ${{}^{s}{\mathcal{U}}}$. By [DD, Lemma 3], ${{}^{s}{\mathcal{U}}}$ is contractible, so ${{}^{s}{\mathcal{U}}}_{E}$ is also contractible. It follows that the product ${{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}$ is contractible. This completes the proof of the lemma. $\square$
End of the proof of Theorem 4.1.   Lemmas 4.2 and 4.3 together with the reasoning in the beginning of the proof show that every adapted trivialization of $\mathcal{H}$ over a closed subspace $Y\subset X$ can be extended to a global adapted trivialization. If $\tau_{0}$ and $\tau_{1}$ are trivializations of $\mathcal{H}$ adapted to $\Pi$, then $\tau_{0}\sqcup\tau_{1}$ is a trivialization of $\tilde{\mathcal{H}}$ over $X\times\left\{0,1\right\}$ adapted to $\tilde{\Pi}$, where $\tilde{\mathcal{H}}$ and $\tilde{\Pi}$ are the liftings of $\mathcal{H}$ and $\Pi$ to $X\times[0,1]$. The polarization $\tilde{\Pi}$ is tame, so $\tau_{0}\sqcup\tau_{1}$ can be extended from $X\times\left\{0,1\right\}$ to a trivialization of $\tilde{\mathcal{H}}$ adapted to $\tilde{\Pi}$. This provides a homotopy between $\tau_{0}$ and $\tau_{1}$ in the class of trivializations adapted to $\Pi$ and completes the proof of the theorem. $\square$
Finite covering dimension.   The proof of Theorem E in the next section is based on the following result.
Theorem 4.4.
Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a metrizable space $X$ of locally finite covering dimension. Then every polarization $\Pi$ of $\mathcal{H}$ is tame and thus satisfies all the conclusions of Theorem 4.1.
Proof. Every subspace $X^{\prime}$ of $X$ is metrizable and thus paracompact, and has locally finite covering dimension by [E, Theorem 7.3.1]. By [DD, Theorem 5], every separable Hilbert bundle over $X^{\prime}$ of infinite rank is locally trivial. The bundle $\mathcal{H}$ is trivial by [DD, Theorem 1] and has separable fiber, so it is separable in the sense of [DD], and so are its subbundles over $X^{\prime}$. Therefore, every infinite rank subbundle of $\mathcal{H}$ over $X^{\prime}$ is locally trivial.
Let $(X_{i},\mathcal{E}_{i})$ be an atlas of $\Pi$. By the reasoning above, both $\mathcal{E}_{i}$ and $\mathcal{E}_{i}^{\bot}$ are locally trivial. The direct sum of their local trivializations provides a local trivialization of $\mathcal{H}$ which takes $\mathcal{E}_{i}$ to a constant subspace and thus is adapted to $\mathcal{E}_{i}$. Therefore, $\Pi$ is tame. Since $X$ is paracompact, $\Pi$ satisfies the assumptions of Theorem 4.1. $\square$
Remark.
For a general space $X$, such a result is no longer true. Indeed, Dixmier and Douady constructed a subbundle $\mathcal{E}$ of the trivial Hilbert bundle $\mathcal{H}$ over a compact space $X$, which is not locally trivial [DD, Theorem 6]. We recall their construction in Example 5.5 below. Every fiber of $\mathcal{E}$ has infinite rank and corank, so $\mathcal{E}$ determines a polarization of $\mathcal{H}$. However, $\mathcal{E}$ has no adapted trivializations, even locally.
5 Self-adjoint operators
Adapted trivializations.   We say that a (local or global) trivialization of $\mathcal{H}$ is adapted to a family $\mathcal{A}$ of regular operators $\mathcal{A}_{x}\in\mathcal{R}(\mathcal{H}_{x})$ if it takes $\mathcal{A}$ to a Riesz continuous family of operators in $H$. If operators $\mathcal{A}_{x}$ are bounded (say, they are projections), then an adapted trivialization takes $\mathcal{A}$ to a norm continuous family. So the definition from the previous section, of a trivialization adapted to a subbundle, agrees with the present definition.
Bounded operators.   Let $\mathcal{A}$ be a family of self-adjoint bounded operators acting on fibers of a Hilbert bundle $\mathcal{H}$. We want to know whether $\mathcal{A}$ admits a global adapted trivialization. Clearly, for that to be true, $\mathcal{A}$ has to be norm quasi-continuous (that is, admit local adapted trivializations). The following theorem shows that for essentially unitary operators this local condition is also sufficient. The author does not know whether this is true for arbitrary bounded self-adjoint operators.
Theorem 5.1.
Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a paracompact space $X$ and $\mathcal{A}$ be a family of self-adjoint essentially unitary operators. Suppose that $\mathcal{A}$ is norm quasi-continuous. Then there exists a global trivialization of $\mathcal{H}$ adapted to $\mathcal{A}$. Moreover, every adapted trivialization of $\mathcal{H}$ over a closed subset $Y$ of $X$ can be extended to a global adapted trivialization. In particular, every two such trivializations are homotopic (in the space of trivializations adapted to $\mathcal{A}$).
Proof. Let $X^{+}$, resp. $X^{-}$ be the subset of points $x\in X$ for which $\mathcal{A}_{x}$ is essentially positive, resp. negative, and let $X^{\star}=X\setminus(X^{+}\cup X^{-})$. Since $\mathcal{A}$ is norm quasi-continuous, each of the three subspaces $X^{+}$, $X^{-}$, and $X^{\star}$ is open and thus closed in $X$, so it is enough to prove the theorem for each of these subspaces separately.
For $x\in X^{+}$, the operator $\mathcal{A}_{x}$ is essentially positive and essentially unitary, so $\mathcal{A}_{x}-1$ is compact. Over $X^{+}$ the family $x\mapsto\mathcal{A}_{x}-1$ is a norm quasi-continuous family of compact operators, so both $\mathcal{A}-1$ and $\mathcal{A}$ are norm continuous in every trivialization of $\mathcal{H}$ over $X^{+}$. The similar reasoning shows that $\mathcal{A}$ is norm continuous in every trivialization of $\mathcal{H}$ over $X^{-}$. Trivializations of $\mathcal{H}$ over $X^{+}\cup X^{-}$ are identified with maps $X^{+}\cup X^{-}\to{{}^{s}{\mathcal{U}}}(H)$, so every two trivializations are homotopic.
It remains to prove the theorem over $X^{\star}$. Without loss of generality we can suppose that $X^{\star}=X$. Consider the family of spectral projections
(5.1) $p_{\lambda}(x)=\operatorname{\mathbb{1}}_{[\lambda,+\infty)}(\mathcal{A}_{x})\;\;\text{for}\;\;x\in X_{\lambda}=\left\{x\in X\,|\,\lambda\in\operatorname{Res}(\mathcal{A}_{x})\right\}.$
The sets $X_{\lambda}$ with $\lambda$ running the interval $(-1,1)$ form an open covering of $X$. A local trivialization of $\mathcal{H}$ adapted to $\mathcal{A}$ takes $p_{\lambda}$ to a norm continuous family over $X_{\lambda}$. For $\lambda<\mu$ the difference
(5.2) $p_{\lambda}(x)-p_{\mu}(x)=\operatorname{\mathbb{1}}_{[\lambda,\mu)}(\mathcal{A}_{x})=\operatorname{\mathbb{1}}_{[\lambda,\mu]}(\mathcal{A}_{x})$
is a norm continuous family of finite rank operators over $X_{\lambda}\cap X_{\mu}$. Hence the collection $(X_{\lambda},p_{\lambda})$, $\lambda\in(-1,1)$ is an atlas of a tame polarization $\Pi_{\mathcal{A}}$ of $\mathcal{H}$ over $X$. The difference $k_{\lambda}=\mathcal{A}-(2p_{\lambda}-1)$ is a norm continuous family of compact operators over $X_{\lambda}$ in a local trivialization adapted to $\mathcal{A}$, so it is norm continuous in every trivialization of $\mathcal{H}$ over $X_{\lambda}$. A trivialization of $\mathcal{H}$ over $X_{\lambda}$ is adapted to $\mathcal{A}=k_{\lambda}+(2p_{\lambda}-1)$ if and only if it is adapted to $p_{\lambda}$. Hence a trivialization is adapted to $\mathcal{A}$ if and only if it is adapted to $\Pi_{\mathcal{A}}$. It remains to apply Theorem 4.1. $\square$
Remark.
In fact, the main part of Theorem 5.1, about operators which are neither essentially positive nor essentially negative, is equivalent to Theorem 4.1. Indeed, let $(X_{i},p_{i})$ be an atlas of a polarization $\Pi$. We associate with it the family $\alpha$ of self-adjoint essentially unitary operators $\alpha(x)=\sum_{i}\rho_{i}(x)r_{i}(x)$, where $(\rho_{i})$ is a partition of unity subordinated to the covering $(X_{i})$ and $r_{i}=2p_{i}-1$. Then $\alpha-r_{i}=\sum_{j}\rho_{j}(r_{j}-r_{i})$ is a norm continuous family of compact operators over $X_{i}$. Therefore, a (local or global) trivialization of $\mathcal{H}$ is adapted to $\Pi$ if and only if it is adapted to $\alpha$.
Proof of Theorem D.   The bounded transform takes self-adjoint operators with compact resolvents to self-adjoint essentially unitary operators and Riesz (quasi-)continuous families to norm (quasi-)continuous families. Therefore, Theorem D follows immediately from Theorem 5.1.
Spectral projections.   For a family $\mathcal{A}_{x}$, $x\in X$ of self-adjoint operators and $\lambda\in\mathbb{R}$ we consider the family $p_{\lambda}$ of spectral projections over $X_{\lambda}\subset X$ defined by (5.1).
Proposition 5.2.
If $\mathcal{A}\colon X\to\mathcal{R}^{\mathrm{sa}}(H)$ is graph continuous, then $p_{\lambda}\colon X_{\lambda}\to\mathcal{P}(H)$ is strong continuous for every $\lambda\in\mathbb{R}$.
Proof. Without loss of generality we can suppose that $X={{}^{g}\mathcal{R}}^{\mathrm{sa}}(H)$ and $\mathcal{A}$ is the identity map. In other words, we need to prove that the spectral projection
$p_{\lambda}=\operatorname{\mathbb{1}}_{[\lambda,+\infty)}\colon\mathcal{R}^{\mathrm{sa}}_{\lambda}\to\mathcal{P}(H),\quad\mathcal{R}^{\mathrm{sa}}_{\lambda}=\left\{A\in\mathcal{R}^{\mathrm{sa}}(H)\,|\,\lambda\in\operatorname{Res}(A)\right\}$
[…]
Proposition 5.3.
Let $\mathcal{A}\colon X\to\mathcal{R}^{\mathrm{sa}}(H)$ be graph continuous and $\lambda\in\mathbb{R}$. Then $\mathcal{A}$ is Riesz continuous on $X_{\lambda}$ if and only if $p_{\lambda}\colon X_{\lambda}\to\mathcal{P}(H)$ is norm continuous.
Proof. Without loss of generality we can suppose that $X_{\lambda}=X$, that is, $\lambda$ belongs to the resolvent set of each $\mathcal{A}_{x}$, $x\in X$.
Suppose that $\mathcal{A}$ is Riesz continuous. Then the family $a_{x}=\bm{\chi}(\mathcal{A}_{x})$ is norm continuous and $\mu=\bm{\chi}(\lambda)$ lies in the resolvent set of each $a_{x}$. Therefore, the map $x\mapsto p_{\lambda}(x)=\operatorname{\mathbb{1}}_{[\mu,+\infty)}(a_{x})$ is also norm continuous.
Conversely, suppose that $p_{\lambda}$ is norm continuous. Then $r_{\lambda}=2p_{\lambda}-1$ is a norm continuous family of symmetries commuting with $\mathcal{A}$. The family $\mathcal{A}^{\prime}=\mathcal{A}\cdot r_{\lambda}$ of self-adjoint operators is graph (resp Riesz) continuous if and only if $\mathcal{A}$ is graph (resp Riesz) continuous. But, in contrast with $\mathcal{A}$, the family $\mathcal{A}^{\prime}$ is uniformly bounded from below by $-|\lambda|$. The graph and Riesz topology coincide on the set of operators bounded from below by a given constant [$P_{2}$, Proposition 11.3], so $\mathcal{A}^{\prime}$ and thus also $\mathcal{A}$ are Riesz continuous. $\square$
Polarization associated with operator family.   Let $\mathcal{A}$ be a graph continuous family of self-adjoint operators with compact resolvents acting on fibers of a Hilbert bundle $\mathcal{H}$. Then $X_{\lambda}$ is open for every $\lambda\in\mathbb{R}$ by [Ka, Theorem IV.3.1] and these sets form an open covering of $X$. By Proposition 5.2, $p_{\lambda}$ is strongly continuous in local trivializations of $\mathcal{H}$, so it determines a decomposition of $\mathcal{H}$ over $X_{\lambda}$ into the orthogonal sum $\mathcal{H}_{\lambda}^{+}\oplus\mathcal{H}_{\lambda}^{-}$ of two Hilbert subbundles. For $\lambda<\mu$ the difference (5.2) is a norm continuous family of finite rank operators over $X_{\lambda}\cap X_{\mu}$ by [Ka, Theorem IV.3.16].
Suppose additionally that each $\mathcal{A}_{x}$ is neither essentially positive nor essentially negative. Then $p_{\lambda}(x)\in\mathcal{P}^{\star}(\mathcal{H}_{x})$, so the collection $(X_{\lambda},p_{\lambda})$ is an atlas of a polarization. Hence a family $\mathcal{A}$ satisfying assumptions above determines a polarization of $\mathcal{H}$ which we call the polarization associated with $\mathcal{A}$ and denote by $\Pi_{\mathcal{A}}$. (The associated polarization $\Pi_{\mathcal{A}}$ is actually finite, but we are interested only in its equivalence class with respect to compact deformations, so we will consider it as a compact polarization.)
Proposition 5.4.
Let $\mathcal{A}$ be a graph continuous family of operators $\mathcal{A}_{x}\in\mathcal{R}^{\star}_{K}(\mathcal{H}_{x})$ acting on fibers of a Hilbert bundle $\mathcal{H}$. A trivialization of $\mathcal{H}$ is adapted to $\mathcal{A}$ if and only if it is adapted to the associated polarization $\Pi_{\mathcal{A}}$.
Proof. This follows immediately from Proposition 5.3. $\square$
Remark.
More generally, consider a family $\mathcal{A}$ of Fredholm self-adjoint operators acting on fibers of $\mathcal{H}$. Suppose that $\mathcal{A}$ is graph quasi-continuous (that is, $\mathcal{A}$ admits local trivializations taking it to a graph continuous family). Then the associated polarization $\Pi_{\mathcal{A}}$ is defined in a similar manner. The only change is in the definition of the set $X_{\lambda}$: one should add the condition that the interval $[\lambda,0]$ or $[0,\lambda]$, depending on the sign of $\lambda$, is disjoint from the essential spectrum of $\mathcal{A}_{x}$.
If a trivialization is adapted to such a family $\mathcal{A}$, then it is adapted to $\Pi_{\mathcal{A}}$ as well. However, the second direction of Proposition 5.4 is no longer true: a trivialization adapted to $\Pi_{\mathcal{A}}$ is not necessarily adapted to $\mathcal{A}$, because it might turn $\mathcal{A}$ into a graph discontinuous family and then Proposition 5.3 cannot be applied. This is the place where condition of compact resolvents comes into the picture: the graph continuity of a family of operators with compact resolvents is independent of the choice of a trivialization.
Proof of Theorem E.   Let $\Pi_{\mathcal{A}}$ be the polarization of $\mathcal{H}$ associated with $\mathcal{A}$. By Proposition 5.4, a (local or global) trivialization of $\mathcal{H}$ is adapted to $\mathcal{A}$ if and only if it is adapted to $\Pi_{\mathcal{A}}$. It remains to apply Theorem 4.4. $\square$
Example 5.5.
Let us show that Theorem E does not hold without additional assumptions on the base space $X$. We will construct a graph continuous family $\mathcal{A}\colon X\to\mathcal{R}^{\star}_{K}(H)$ of invertible self-adjoint operators with compact resolvents over a compact metrizable space $X$ such that $\mathcal{A}$ is not Riesz quasi-continuous. Moreover, for every point $x\in X$, $\mathcal{A}$ is not Riesz quasi-continuous at $x$ (that is, it is not Riesz quasi-continuous in any neighborhood of $x$).
Our example is based on a construction of a non-trivial separable Hilbert bundle $\mathcal{E}$ over a compact space $X$ by Dixmier and Douady. See [DD, Theorem 6]. Let us recall their construction. Let $H_{i}=\mathbb{C}^{2}$, $X_{i}=\mathbb{P}(H_{i})\cong\mathbb{C}\mathbb{P}^{1}$, $H=\oplus_{i=1}^{\infty}H_{i}$, and $X=\prod_{i=1}^{\infty}X_{i}$ with the product topology. The space $X$ is a countable product of metrizable spaces and thus is metrizable itself; it is a product of compact spaces and thus is compact itself. The Hilbert bundle $\mathcal{E}$ over $X$ is defined as the subbundle of the trivial Hilbert bundle $H_{X}$ with the fiber $\mathcal{E}_{x}=\oplus_{i=1}^{\infty}x_{i}$ over $x=(x_{i})$, where $x_{i}\in X_{i}$ is considered as a one-dimensional subspace of $H_{i}$. By [DD, Theorem 6], every section of $\mathcal{E}$ vanishes at some point. In particular, $\mathcal{E}$ is not trivial and thus it is not locally trivial (since all locally trivial Hilbert bundles over paracompact base spaces are trivial).
Moreover, the bundle $\mathcal{E}$ looks the same near each point of $X$ (more precisely, the group $\prod_{i=1}^{\infty}\mathcal{U}(H_{i})$ acts transitively on $X$ and this action lifts to the action on $\mathcal{E}$ by bundle automorphisms), so a local triviality of $\mathcal{E}$ near one point would imply local triviality over the whole $X$. Therefore, for every open subset $X^{\prime}\subset X$, $\mathcal{E}$ is not trivial over $X^{\prime}$.
Let $c\colon\mathbb{N}\to\mathbb{R}_{>0}$ be a sequence of positive numbers converging to $\infty$. Then
$A=\oplus_{i=1}^{\infty}c_{i}\cdot\operatorname{Id}_{H_{i}}$
is a positive regular operator with compact resolvents. Let $p(x)=\oplus_{i=1}^{\infty}p_{i}(x)$ be the orthogonal projection of $H$ onto $\mathcal{E}_{x}$, and let $r(x)=2p(x)-1$ be the corresponding symmetry. Then $r(x)$ commutes with $A$, and $\mathcal{A}(x)=A\cdot r(x)$ is an invertible self-adjoint operator with compact resolvents, which is neither essentially positive nor essentially negative.
We claim that the family $\mathcal{A}\colon X\to\mathcal{R}^{\star}_{K}(H)$ is graph continuous. Indeed, $(1+\mathcal{A}_{x}^{2})^{-1}=(1+A^{2})^{-1}$ is a compact operator independent of $x$. The map
$x\longmapsto\mathcal{A}_{x}(1+\mathcal{A}_{x}^{2})^{-1}=r_{x}A(1+A^{2})^{-1}$
is the product of a strongly continuous family of symmetries $r_{x}$ and a fixed compact operator $A(1+A^{2})^{-1}=\bm{\chi}(A)(1+A^{2})^{-1/2}$ and thus is norm continuous. Therefore, all components of the orthogonal projection (2.6) onto the graph of $\mathcal{A}_{x}$ are norm continuous and thus $\mathcal{A}$ is graph continuous.
However, for every open subset $X^{\prime}\subset X$, the restriction of $\mathcal{A}$ to $X^{\prime}$ is not Riesz quasi-continuous. Indeed, the positive spectral projection $p(x)$ of $\mathcal{A}(x)$ is the orthogonal projection of $H$ onto $\mathcal{E}_{x}$. If $\mathcal{A}$ were Riesz continuous in some local trivialization of the trivial Hilbert bundle $X^{\prime}\times H$, then the map $x\mapsto p(x)$ would be norm continuous in the same local trivialization and $\mathcal{E}$ would be a locally trivial Hilbert bundle over $X^{\prime}$, which is not the case.
Remark.
If one multiplies the family $\mathcal{A}_{x}$ from this example by the strongly continuous family of unitary operators $r_{x}$, then the resulting family $x\mapsto\mathcal{A}_{x}\cdot r_{x}=A$ will be Riesz continuous (and even constant). However, for every strongly continuous family of unitary operators $v_{x}\in\mathcal{U}(H)$, the family $x\mapsto v_{x}\mathcal{A}_{x}v_{x}^{*}$ is Riesz discontinuous at every point $x\in X$. This demonstrates the difference between Theorems C and D, that is, between one-sided multiplication and conjugation by unitary operators, which we discussed in the Introduction.